### v1

# mO Haoyu[#1]

Abstract
Segment Anything Model (SAM) has shown impressive performance in interactive and zero-shot segmentation across diverse domains, suggesting that they have learned a general concept of “objects” from their large-scale training. However, we observed that SAM struggles with certain types of objects, particularly those featuring dense, tree-like structures and low textural contrast from their surroundings. These failure modes are critical for understanding its limitations in real-world use. In order to systematically examine this issue, we propose metrics to quantify two key object characteristics: tree-likeness and textural separability. Through extensive controlled synthetic experiments and testing on real datasets, we demonstrate that SAM’s performance is noticeably correlated with these factors. We link these behaviors under the concept of “textural confusion”, where SAM misinterprets local structure as global texture, leading to over-segmentation, or struggles to differentiate objects from similarly textured backgrounds. These findings offer the first quantitative framework to model SAM’s challenges, providing valuable insights into its limitations and guiding future improvements for vision foundation models.
Introduction
Segment Anything Model (SAM) [27], the most widely-used foundation model for image segmentation, has demonstrated promising zero-shot and fine-tuned segmentation ability for objects not seen in training, within a wide range of domains since its release, such as biomedical images [30, 22, 10], remote sensing [6, 35], etc. However, certain failure modes for SAM have been found empirically, where SAM produces surprisingly underwhelming performance on datasets with specific uncommon objects or atypical contexts, such as those with dense, branching structures (e.g., retinal blood vessels) [36, 11], concealed/low-contrast objects [7, 39, 24], small or irregular objects, and others [24] [^1] We focus on “unintentional” failure modes of SAM, as opposed to intentional adversarial attacks/examples [45, 47].. A solution to this problem could simply be fine-tuning the model on such challenging examples using various strategies [29, 32, 17], but this would still lack an understanding of the underlying causes of the issues, and so may not result in the best solution. As such, it is important to develop a better theory for why these failures happen, in order to better understand (and potentially mitigate) the limitations of SAM for new applications.
Figure 1: SAM’s segmentation performance tends to drop noticeably when the object has high tree-likeness (left, on DIS) or low textural separability (right, on iShape)–even with heavy prompting–which we investigate in this work. From left to right: input images, object annotations, and their SAM segmentations.
In this paper, we propose that several of these failure modes can be explained as SAM misinterpreting textural cues for how it disambiguates an object from its surroundings: it typically confuses objects with irregular, dense tree-like local structure as large-scale textures (resulting in over-segmentation), and similarly is generally challenged by objects that have similar textures to their surroundings. We propose metrics developed from first principles for quantifying the characteristics of (1) object tree-likeness and (2) textual separability. Objects tree-likeness (Sec. 2.2) is based on concepts such as spatial contour frequency and the difference in typical object variability between global and local scales. Textural separability (Sec. 3.1) is based on the detectability of differences in textural features of an object compared to its surroundings. We designed these metrics to be simple and interpretable, such that they can be applied to any dataset or object type. We also present them in quickly-computable PyTorch-like algorithmic form that can be applied to any masks.
Next, we begin to probe the effect that these characteristics have on SAM’s segmentation ability through experiments on synthetic data which are designed to cover a range of object tree-likeness (Sec. 2.3.1) and textural separability (Sec. 3.2.1). These experiments are carefully controlled to disentangle the effects of object shape and texture on performance, as well as mitigate any other potential confounding factors. We then continue with similar experiments on real datasets (Secs. 2.3.2 and 3.2.2) that represent a wide range of objects to segment. In all experiments, we find both the exact measured tree-likeness and the textural separability of objects are noticeably correlated with SAM’s performance (IoU) in segmenting it, over a wide range of object and image types. Our findings are the first to quantitatively model and experimentally verify how the performance of segmentation foundation models is affected by certain measurable object characteristics, providing an explanation for prior findings of SAM’s unexpected failures on new datasets.
Our overall contributions are as follows:
  1. 1.
    Based on first principles, we propose quantitative metrics for object tree-likeness of Contour Pixel Rate (CPR) and Difference of Gini Impurity Deviation (DoGD), as well as a metric for object textural separability.
  2. 2.
    We present carefully-designed experiments for the segmentation of synthetic images which cover a range of tree-likeness and textural separability, where important factors such as object shape and texture are controlled.
  3. 3.
    We show that SAM’s segmentation performance tends to negatively correlate with an object’s tree-likeness, and positively correlate with it’s textural separability, on a variety of experiments on both synthetic and real images.
We hope that this inspires future work in understanding the limitations of vision foundation models. We release easy-to-use code for our proposed metrics at https://github.com/mazurowski-lab/SAM-TexturalConfusion-Metrics.
Related Works
In addition to the various “failure modes” and biases of segmentation foundation models discussed in the introduction which motivated this work, various works have also studied the same for general vision models. Central to these is the goal of understanding the visual features that networks use to make predictions from images, and whether object shape or texture is more important. This is important to study because any unexpected failure of such models will depend on the features used for inference.
Zhang et al. [46] posited that objects within images can be described by three characteristics: shape, texture, and the composition of textures within shapes, and analyzed how these characteristics may affect the challenge of segmenting such objects. It is well known that convolutional neural networks’ (CNNs’) predictions are typically biased towards the texture of objects rather than their shape, for both classification [16, 38, 21, 3] and segmentation [46], resulting in curious examples where network predictions are based on texture even when the shape clearly defines a different object. For example, [28] found that when a CNN is presented with images of objects with only their shape visible, not texture (silhouettes), performance was greatly worsened.
It has also been found that the shape-learning behavior of neural networks can differ noticeably depending on the particular training dataset and objective [21, 20], motivating us to study the specific behavior of SAM (given its’ unique training dataset and pipeline) in detail. Interestingly, there has been some evidence that vision transformers (like SAM’s image encoder) may instead be biased towards shape, rather than texture for classification tasks [40]. Here, we study segmentation rather than classification, where it is unclear if those results extend, as our results generally point to SAM focusing on texture over shape (to the point of even confusing complex, locally-varying shape itself as texture).
1 Methods: SAM Usage and Prompting
We experiment with both the ViT-H (default) and ViT-B standard pretrained SAM models [27]. As is default for SAM, all input images are resized to a resolution of $1024\times 1024$, and normalized to $[0,255]$. All images that possess instance segmentations for multiple objects (e.g., iShape and Plittersdorf) will be evaluated by SAM segmenting each of the images’ objects one at a time (guided by clear prompting), with all other mask pixels set as background. In all experiments, we use the oracle-chosen mask out of SAM’s three outputted predictions.
For all experiments on the synthetic images of Sec. 2.3.1, we provide a tight bounding box about the object of interest as the prompt. For all experiments on real images as well as the style-transferred images of Sec. 3.2.1, we provide the same bounding box as well as a certain number of positive and/or negative point prompts randomly sampled from the object foreground and background within the bounding box, respectively, depending on the dataset (full details in Appendix C.1). We use this relatively heavy prompting strategy in order to minimize the ambiguity of instructions provided to SAM, which will also help to minimize any differences of the oracle prediction from the other two.
2 The Challenge of Tree-Like Structures
2.1 Motivation
We first wished to understand SAM’s challenges in segmenting certain objects with dense, tree-like structures, following our observation of a trend of SAM having difficulty with such objects including retinal blood vessels [34, 29] or satellite road images [14, 42] (see e.g., Fig. 2). Interestingly, while both of these object types have branching features, a characteristic which naively could relate to SAM’s failure, SAM’s performance on retinal vessel images is noticeably worse (avg. $\mathrm{IoU}\simeq 0.05$ [34]) compared to on satellite road structures (avg. $\mathrm{IoU}\simeq 0.2$ [14]), which was reproduced in our own experiments (Fig. 4).
Figure 2: Example retinal blood vessel (top) and satellite road (bottom) images and accompanying object segmentation masks.
This motivated us to quantify the features of these types of objects that relate to SAM’s failure. We hypothesize that such objects become more challenging to segment when their branching structures are dense and irregularly-spaced, which we refer to as “tree-like”. We will show that SAM considers such structures as textures rather than shapes, resulting in significant over-segmentation. In the following section we propose how to characterize the tree-likeness of objects using two new quantitative metrics: CPR (Contour Pixel Rate) and DoGD (Difference of Gini Impurity Deviation).
2.2 Quantifying Tree-Like Structures
2.2.1 Contour Pixel Rate
Consider some image $x\in\mathbb{R}^{C\times H\times W}$ with a “ground truth” binary segmentation mask for some object within it, $m\in\{0,1\}^{H\times W}$ (in general, this could also be just one class of some multi-class segmentation mask). We first propose to measure the degree of tree-like structure of the object $m$ according to the percentage of the object’s pixels which lie on it’s contour, which are defined as follows.
Definition 1 (Contour Pixels).
Given some mask $m$, a pixel $m_{ij}$ is a contour pixel if $m_{ij}=1$ and there exists a different pixel $m_{kl}$ such that $m_{kl}\neq m_{ij}$ and $||(k,l)-(i,j)||_{1}<R$, given some small contour width threshold $R>0$.
In other words, contour pixels have at least one pixel of a different class within a small neighborhood. Taking $F$ to be the set of foreground pixels $F:=\{(i,j):m_{ij}=1\}$, the set of contour pixels of $m$ is then $C=\{(i,j)\in F:\exists(k,l)\in F^{c}$ such that $||(k,l)-(i,j)||_{1}<R\}$, which we use to define the Contour Pixel Rate (CPR) of the object as
$\mathrm{CPR}(m):=\frac{|C|}{|F|}.$ (1)
Intuitively, tree-like objects will have a higher percentage of contour pixels, resulting in higher CPRs. We demonstrate fast computation of CPR in full detail in Algorithm 1, via vectorized PyTorch-like pseudocode.
Algorithm 1 Contour Pixel Rate (CPR) of an object (PyTorch-like pseudocode).
Input: Object mask ($H\times W$ tensor), contour width threshold R (int).
from skimage.morphology import diamond
def CPR(mask, R):
neighb_kernel = diamond(R)
neighb_counts = conv2d(mask,
neighb_kernel, padding=R)
contour_pix = logical_and(mask>0,
neighb_counts <= (neighb_kernel.sum()-1))
cpr = contour_pix.sum()/mask.sum()
return cpr.item()
2.2.2 Difference of Gini-impurity Deviation (DoGD)
We alternatively propose to measure the tree-likeness of some object according to how the variability of object presence across different locations in the image differs between global and local scales. Intuitively, irregularly-spaced tree-like structures have high variability at small scales due to alternating frequently between areas of mixed and uniform pixel classes, but low variability at large scales due to the repetitive nature of the structure becoming more homogeneous, which we propose to quantify as follows.
First, we quantify the object presence within some $k\times k$ square window of the mask anchored at some coordinates $h_{0}<H$, $w_{0}<W$ via the Gini impurity:
$\mathrm{Gini}(m;k,h_{0},w_{0}):=1-{\sum}_{j}[p_{j}(\mathcal{W}^{k}_{h_{0},w_{0}}(m))]^{2}$ (2)
where we denote the $k\times k$ square window of the mask anchored at $h_{0},w_{0}$ as $\mathcal{W}^{k}_{h_{0},w_{0}}(m):=m[h_{0}:h_{0}+k,w_{0}:w_{0}+k]$. Here, $p_{j}(\mathcal{W}^{k}_{h_{0},w_{0}}(m))$ denotes the probability of the object of class $j$ being in some pixel within the window, which is computed simply as $n_{j}/k^{2}$, where $n_{j}$ is the number of pixels in the window of class $j$. In our binary segmentation case, the Gini impurity simplifies to
$\mathrm{Gini}(m;k,h_{0},w_{0}):=$ (3)
$1-[p(\mathcal{W}^{k}_{h_{0},w_{0}}(m))]^{2}-[1-p(\mathcal{W}^{k}_{h_{0},w_{0}}(m))]^{2},$
[…]
Next, we compute the variability of object presence at a given scale/window size across the entire mask by sampling all possible $k\times k$ windows with anchors $(h_{0},w_{0})$, and computing the standard deviation of the Gini impurity across these windows, as
$\sigma^{\mathrm{Gini}}_{k}(m):=\sqrt{\mathrm{Var}_{h_{0},w_{0}}[\mathrm{Gini}(m;k,h_{0},w_{0})]}.$ (4)
Finally, we define the Difference of Gini Impurity Deviation (DoGD) between global and local scales as
$\mathrm{DoGD}(m):=\sigma^{\mathrm{Gini}}_{a}(m)-\sigma^{\mathrm{Gini}}_{b}(m),$ (5)
where the global and local window sizes $k=a$ and $k=b$ are chosen such that $a\gg b$. We present DoGD in optimized PyTorch-like form in Algorithm 2.
Algorithm 2 Difference of Gini Impurity Deviation (DoGD) of an object (PyTorch-like pseudocode).
Input: Object mask ($H\times W$ tensor), global and local window sizes a, b (ints).
def DoGD(mask, a, b):
gini_std = {}
for k in [a,b]:
avg_kernel = ones(k,k)
p = conv2d(mask, avg_kernel)
gini = 1 - p**2. - (1-p)**2.
gini_std[k] = gini.std().item()
return gini_std[a] - gini_std[b]
Intuitively, objects with significant tree-like or fractal-like structure will exhibit relatively large values of $\sigma^{\mathrm{Gini}}_{b}(m)$ due to high variability in pixel composition at small scales (frequently alternating between areas of mixed classes and areas of a single class), yet small $\sigma^{\mathrm{Gini}}_{a}(m)$ due to structures with high uniformity and/or repetitions at large scales, altogether increasing the DoGD.
We perform all experiments with the hyperparameters for CPR and DoGD set to $R=5$, $a=127$, and $b=3$, which we found via grid search for the values which resulted in the Kendall’s $\tau$ between IoU and DoGD with the lowest $p$-value on the held-out DIS training set using ViT-H SAM. We show results using a wide range of other values for these hyperparameters in Appendix B.1, where we found our findings to be consistent for most other settings. Moreover, we note that CPR and DoGD are correlated with each other (Pearson $|r|=0.83$ on average; Appendix D.1), showing their consistency in how they quantify different aspects of tree-likeness.
2.3 The Relationship between Object Tree-likeness and Segmentation Performance
2.3.1 Experiments on Synthetic Data
In this section, we will first carefully probe the effect of object tree-likeness on SAM’s segmentation performance by testing it on synthetic images which solely possess objects of varying tree-likeness, with different independently chosen, uniform foreground and background textures. These objects are contiguous components samples from retinal blood vessel and satellite road masks, with the full procedure of generating these images detailed in Appendix A.2. Example generated images, masks and SAM segmentation predictions for them are shown in Fig. 3. As shown (as well in Fig. 4), the objects cover a wide range of tree-likeness as measured by these quantities.
In order to mitigate any confounding on SAM performance due to an object’s textural contrast from its surroundings (which we study in Sec. 3), for each mask $m_{c}$ generated by our procedure, we apply SAM to $K=7$ images created by applying $K$ different randomly-sampled pairs of textures to the object’s foreground ($m_{c}=1$) and background ($m_{c}=0$). We then obtain SAM’s final prediction for this object via pixel-wise majority voting over its predictions on these $K$ images.
Figure 3: Left: Example synthetic tree-like images and object masks. Right: Trend of increasing tree-likeness (increasing CPR/decreasing DoGD) of these objects resulting in worse SAM segmentation predictions (to the right of each object mask).
In Fig. 4 we show the relationship between an object’s tree-likeness (as measured by CPR or DoGD) with SAM’s performance (IoU) in segmenting the object, for all generated synthetic objects. We quantify the strength of this relation via rank/non-linear correlation, measured by Kendall’s tau ($\tau$) [25] and Spearman’s rho ($\rho$) [37], shown in Table 1333We evaluate $\tau$ in addition to $\rho$ due to it being more robust to outliers.. Intriguingly, we see that object tree-likeness is quite predictive of SAM’s performance, with average absolute correlations of $|\tau|=0.76$ and $|\rho|=0.93$ for CPR and $|\tau|=0.64$ and $|\rho|=0.82$ for DoGD; i.e., more prevalent dense tree-like structures corresponding to worse performance.
CPR DoGD
SAM Encoder $\tau$ $\rho$ $\tau$ $\rho$
ViT-H -0.77 -0.93 0.61 0.80
ViT-B -0.75 -0.93 0.66 0.84
Table 1: Rank correlation between SAM segmentation IoU and object tree-likeness (CPR and DoGD), on the synthetic dataset.
Figure 4: SAM segmentation IoU vs. object tree-likeness (CPR, left; DoGD, right) on the synthetic dataset.
While this relationship is certainly strong, it is still on synthetic, controlled data. In the following section, we will show that this finding is also present for real data which is subject to noise from a variety of uncontrollable factors.
2.3.2 Experiments on Real Data
We will now perform the same analysis on two real datasets which contain objects which cover a spectrum of tree-likeness, DIS5k and iShape. DIS5k [33] (or “DIS” for short) is a dataset that contains extremely detailed segmentation masks of objects with varying degrees of hollowness, and both regular and irregular tree-like structures. All DIS experiments will be reported on its validation set unless otherwise stated. We show example images with objects with varying degrees of tree-likeness (by CPR) in Fig. 1 left.
iShape [43] consists of six sub-datasets of real and realistic-appearing synthetic images for instance segmentation of different objects: antenna, branches, fences, logs, hangers, and wires (see e.g. Fig. 9). We analyze these classes individually to mitigate potential confounding/noise factors due to inter-class variations; we do not do this for DIS due to the larger number of classes which are much more fine-grained in their differences, additionally because there are few images per class for DIS.
Results.
In Fig. 5 we show how SAM’s performance (IoU) on these images relates to the tree-likeness of the objects which it is segmenting, with accompanying quantitative correlation results shown in Table 2. In order to reduce the noise incurred by the large variety of segmented objects in the dataset and nuisance confounding factors in the images, we analyze results with aggregated objects: we cluster groups of $5$ objects/images with similar IoU and tree-likeness (CPR or DoGD) into single datapoints of the average value of these metrics, and similar for the IoU vs. textural separability experiments of Sec. 3.
iShape
DIS antenna branch fence hanger log wire
SAM Enc. $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
CPR: ViT-H -0.59 -0.76 -0.13 -0.21 -0.61 -0.81 -0.61 -0.81 -0.25 -0.37 -0.60 -0.77 -0.06 -0.10
ViT-B -0.63 -0.81 -0.18 -0.27 -0.63 -0.82 -0.63 -0.83 -0.44 -0.62 -0.62 -0.80 -0.12 -0.17
DoGD: ViT-H 0.58 0.77 0.23 0.30 0.52 0.70 0.46 0.65 0.52 0.72 0.30 0.46 0.08 0.11
ViT-B 0.45 0.64 0.12 0.14 0.51 0.69 0.45 0.64 0.50 0.70 0.13 0.22 0.06 0.08
Table 2: Rank correlation between SAM prediction IoU and object tree-likeness (CPR, upper table) and DoGD, lower table), on DIS and iShape.
Figure 5: SAM segmentation IoU vs. object tree-likeness (CPR, top; DoGD, bottom) on DIS (left) and iShape (right).
We see that despite the many potential confounding factors in real data, there is still a clear correlation between object tree-likeness as measured by the proposed metrics and SAM segmentation performance. In particular, we see an average absolute correlations of $|\tau|=0.56$ and $|\rho|=0.74$ for CPR and $|\tau|=0.44$ and $|\rho|=0.62$ for DoGD, excluding the antenna and wire objects of iShape which had outlying correlations, likely because those two object classes cover only small ranges of tree-likeness as opposed to the other iShape classes (Fig. 5), such that noise from other confounding factors obscures any dependence of performance on tree-likeness.
3 The Challenge of Textural Separability
In the previous section, we demonstrated that SAM struggles with segmenting non-conventional shapes, in particular, dense tree-like structures, which we hypothesize is due to the model confusing the dense structure as the texture of a non-treelike, more regular shape, rather than a shape itself (see e.g. Fig. 1 left). Similar to this behavior is that even for objects with simpler shapes, SAM can still be confused if the object’s texture is even somewhat similar to its surroundings, which we will study in this section.
3.1 Measuring Textural Separability
We will define the textural contrast or separability between some object mask and its surroundings by how easily their textures can be distinguished from one another. Motivated by findings that early layers of classification-pretrained CNNs primarily capture low-level features involving edges and textures [44], we will characterize an image’s textures using the first convolutional layer of a ResNet-18 [19] pretrained on ImageNet [9]. Denoted by $f_{1}:\mathbb{R}^{C\times H\times W}\rightarrow\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$, this outputs a textural feature map corresponding to all $7\times 7$ windows in the image.
We then measure the textural separability of an object according to if a simple classifier $g$ can be trained to discriminate between (a) the activations for the pixels of the object foreground and (b) the activations right outside of the object boundary [^4] This procedure is somewhat similar to the probing of hidden activation concepts [2, 26], although these works detected concepts (such as textures) at the image level, not at the single activation level as we do here.. We define this explicitly in Algorithm 3, where $\mathrm{Dilate}$ and $\mathrm{Disk}$ refer to scikit-image.morphology functions [41]. We use simple logistic regression for $g$ with inverse regularization parameter $C=2$ [31]. We also evaluate using other hyperparameter settings, as well as a random forest classifier for $g$ instead, in Appendix B.2, where we found similar results.
Algorithm 3 Textural separability of an object.
0:  Image $x\in\mathbb{R}^{C\times H\times W}$ with mask $m\in\{0,1\}^{H\times W}$, first convolutional layer of pretrained CNN $f_{1}$, simple classifier $g$.
1:  $h=f_{1}(x)\in\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$ (Get textural features.)
2:  $m=\mathrm{Resize}(m)$ to $H^{\prime}\times W^{\prime}$ via NN interpolation.
3:  $m^{\prime}:=\mathrm{Dilate}(m,\mathrm{Disk}(5))$ (Expand boundary.)
4:  $m^{\prime}=m^{\prime}-m$ (Extract boundary only.)
5:  $h_{\mathrm{obj}}:=h[:,m\neq 0]$
6:  $h_{\mathrm{bdry}}:=h[:,m^{\prime}\neq 0]$
7:  Train $g$ to classify between activations in $h_{\mathrm{obj}}$ vs. $h_{\mathrm{bdry}}$.
8:  return  Training classification accuracy of $g$.
3.2 The Effect of Textural Separability on Segmentation Performance
3.2.1 Style Transfer for Controlled Textural Contrast
Similar to Sec. 2.3, in studying the effects of textural separability on SAM’s performance, we also wish to perform experiments on well-controlled synthetic images in order to more carefully disentangle the effects of textural separability shape on performance from the effects due to object shape, before evaluating on real data (in the following section). Here, we will do so using neural style transfer [15] to precisely modify the textural contrast of real objects/images. This will involve (1) modifying objects’ shape without altering their background using an inpainting model, followed by (2) applying neural style transfer (NST) to the composite image to adjust the textural contrast of the object with the surrounding image.
We begin with images sampled from the VOC2012 dataset [13] accompanied by objects (instance masks), according to certain criteria. Namely, we pick objects which take up between $5\%$ and $25\%$ of the entire image’s area (to have large enough objects while still maintaining sufficient background), sampling 484 objects total (with accompanying background) for study. For each object, we use a Stable Diffusion-based inpainting model (details in Appendix C.2) to remove the object from its corresponding image and fill in it’s mask area with the background. This creates a pair of a background image and a separate object (mask), allowing us to have careful control of modifying the object (and its mask accordingly) before placing it back into the inpainted background. For a given object, we then create three types of composite images with varying degrees of changes to the object’s shape (without modifying the background):
  1. 1.
    Controlled: the object is not modified.
  2. 2.
    Altered: the object undergoes major non-affine geometric transformations (see App. C.2), resulting in shape modification and its texture being squeezed or expanded.
  3. 3.
    Mixed: this variant uses pixels from the altered image where the original and altered masks overlap, and pixels from the original object where they do not overlap.
After one of these three composite image types is created, we use neural style transfer (NST) to apply a texture to the image (similar to experiments in [16]) randomly sampled from Colored Brodatz [1], via an implementation of the NST model of Gatys et al. [15] (details in Appendix C.2). We illustrate an example of this entire procedure in Fig. 6.
Figure 6: Example images from each component of the synthetic textural separability dataset creation pipeline.
We adjust textural separability by performing the style transfer at eight different degrees of severity using different settings for the weighting hyperparameters that control the balance between content preservation and style transfer (details in Appendix C.2), with example styled images shown in Fig. 7 (more in Appendix. A.3). In Fig. 8 left, we validate this scheme by showing that indeed, steadily increasing the NST intensity results in decreasing textural separability.
Figure 7: Using style transfer with different intensities to adjust the textural separability of an object (highlighted in red), on controlled (upper row) and altered (lower row) versions of the object.
Figure 8: Left: Increasing NST intensity results in decreased textural separability. Right: SAM segmentation IoU vs. NST intensity, on the three composite image types.
In Fig. 8 right, we show how SAM’s segmentation performance changes with respect to style transfer intensity on the test set, for each of the three types of object transformations. We first see that SAM’s performance decreases with lower textural separability/higher NST intensity. Second, there is an interesting pattern of clear gaps in segmentation performance between the three object transformation types. SAM performs best on the altered objects/foregrounds, which have both distorted shape and texture, compared to the un-altered/controlled objects, with the mixed objects in-between. This further demonstrates the effect of textural contrast on segmentation performance, as the altered objects have additional textural contrast to their backgrounds due to the transformations which were applied to their shape, boundary and texture.
3.2.2 Experiments on Real Data
We will now evaluate datasets of real images; iShape (Sec. 2.3.2) and Plittersdorf [18], both of which were used in the original SAM paper [27], and possess objects with a wide range of textural separability from their surroundings. Plittersdorf consists of camera trap video frames of wild deer recorded in a wildlife park. These frames often have low contrast objects due to frequent low-light conditions, making it a useful dataset for this analysis. Example images and objects from both datasets are shown in Fig. 9.
Figure 9: Example images and object masks from Plittersdorf (left group) and iShape (right group).
In Fig. 10, we show how the textural separability (Algorithm 3) of these objects relates to SAM’s segmentation performance on them, with correlation results shown in Table 3. Overall, across all datasets we see a fairly strong correlation between textural separability and segmentation performance (average correlation of $\tau=0.49$ and $\rho=0.66$), especially considering the variety of objects and backgrounds, with objects with low separability resulting in especially poor segmentation ($\mathrm{IoU}\sim 0.3$). In these cases, SAM was confused by objects close to the object of interest which also have similar textures (see e.g., Fig. 10, right).
iShape
antenna branch fence hanger log wire Plittersdorf
SAM Enc. $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
ViT-H 0.44 0.59 0.50 0.68 0.65 0.85 0.63 0.83 0.33 0.48 0.49 0.67 0.26 0.38
ViT-B 0.46 0.61 0.56 0.75 0.67 0.86 0.65 0.84 0.36 0.52 0.55 0.73 0.36 0.50
Table 3: Rank correlation between SAM prediction IoU and object textural separability, on iShape and Plittersdorf.
Figure 10: SAM prediction IoU vs. object textural separability (Algorithm 3), for iShape (left) and Plittersdorf (right).
Discussion
Understanding and Interpreting SAM’s Failure Modes.
First, we see that SAM’s performance issues on cases of high tree-likeness (Sec. 2.3) were primarily due to over-segmentation, rather than under-segmentation; the model would frequently generate large false-positive regions in the empty space of densely packed, thin tree-like structures (e.g., Figs. 3, right and 1, left). This could not simply be due to image resolution/downsampling blurring the thin objects, as for structures which are also thin but more sparsely packed, the over-segmentation is reduced (see e.g., Fig. 3, center rightmost). We therefore hypothesize that the root of SAM’s failure on these objects may lie in the highly repetitive yet irregular, dense patterns inherent to tree-like structures in particular: SAM confuses these patterns of shape as the texture of a more regularly-shaped object.
In the cases of both tree-like and low-textural contrast objects, SAM either has trouble with correctly delineating an object’s shape when faced with a “confusing” textural cue, be it either a dense, irregular shape which appears to instead be a texture, or an object texture that is similar to its surroundings. One explanation for this is that SAM’s training set SA-1B [27] possessed few objects with these qualities. This seems likely for tree-like objects, given the typically low concavity of objects in SA-1B (Fig. 6, right in [27]), which were captured in photographic contexts where such objects are uncommon. Low-textural contrast objects are similarly uncommon in most photographs, and furthermore, segmenting objects is generally more challenging if they are harder to pick out from their surroundings.
Applications of Our Findings and Metrics.
A simple application of our findings and proposed metrics for the tree-likeness and textural separability of objects would be to predict if SAM is expected to perform underwhelmingly on new data, according to if the objects’ measured tree-likeness is high or the textural separability is low. Similarly, the general diversity of objects in a segmentation foundation model’s training/fine-tuning dataset could be measured according to the dataset’s distribution of these metrics, which could inform whether additional training data is needed to be acquired/annotated to develop a better generalist model.
Limitations and Future Work.
For our findings of the correlations between the object characteristics measured by our metrics and SAM’s performance in segmenting such objects, we attempted to minimize the influence of other confounding factors which could effect segmentation performance by first establishing a baseline with our carefully-controlled synthetic data experiments. We then chose real evaluation datasets which covered a wide range of object types according to these metrics in order to gain a better “signal-to-noise ratio” for the studied trends. However, it is impossible to mitigate all potential confounding factors which could affect performance in real data, which is why our correlation findings are still noticeable for the real datasets, yet not quite as tight as on the synthetic data.
We do not evaluate the 2D version of the recently released SAM 2 model due to it likely resulting in similar behavior as SAM 1, along with evidence that oracle predictions of SAM and SAM 2 are typically similar in performance to within $\lesssim 0.1$ IoU [12].
Another future direction would be to explore the specific component(s) of SAM which can be most attributed to these failures (in a mechanistic sense), which could guide further model development. We deem it likely that the susceptible component of SAM is the image encoder, simply because it is the main backbone of the model which extracts features to be used by the lightweight mask decoder, but obtaining a more fine-grained answer would require a careful treatment.
Conclusion
In this paper, we quantitatively modeled how the segmentation performance of SAM relates to certain measurable object characteristics: tree-likeness and textural separability. We find SAM’s performance to typically be noticeably correlated with these factors, showing the need for further work in understanding and potentially mitigating such “failure modes” of vision foundation models.
Acknowledgements
Research reported in this publication was supported by the National Heart, Lung, and Blood Institute of the National Institutes of Health under Award Number R44HL152825. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
Supplementary Material
Appendix A Additional Dataset Details
A.1 Retinal Blood Vessel and Road Satellite Images
The retinal blood vessel and road satellite image object masks which we use to generate the synthetic images for the experiments in Sec. 2.3.1, are from the Retina Blood Vessel [23] and the Road Extraction Challenge data of DeepGlobe18 [8]. For our retinal vessel mask set, we use the vessel masks from the training set of 80 image/mask pairs, and for our road mask set, we randomly sample the same number of masks from the split-merged full dataset, in order to ensure that these two object types appear equally in the synthetic dataset of Sec. 2.3.1.
A.2 Synthetic Tree-like Object Images
Our algorithm for creating synthetic images of tree-like objects (used in Sec. 2.3.1) is shown as follows.
  1. 1.
    Sample object mask $m$ from either retinal blood vessel or satellite road image datasets (details in Appendix A.1).
  2. 2.
    Randomly select one of the contiguous components of $m$, denoted $m_{c}$, and enclose it with a tight bounding box.
  3. 3.
    Resize $m_{c}$ such that its bounding box is $512\times 512$, and randomly place it in a blank $1024\times 1024$ image.
  4. 4.
    Apply two different textures randomly sampled from Colored Brodatz [1] to the foreground ($m_{c}=1$) and background ($m_{c}=0$), resulting in the final image.
In Fig. 12 we show example images generated with different foreground and background texture combinations for various objects.
A.3 Neural Style Transfer (NST) Generated Images
In Fig. 11 we provide additional example images generated via inpainting and neural style transfer for the textural separability experiments of Sec. 3.2.1.
Figure 11: Additional examples of inpainting+neural style transfer-generated images to supplement Fig. 6.
Figure 12: Examples of synthetic tree-like object images. Object masks shown in the right-most column, with images created using randomly chosen different foreground and background texture combinations in the left columns (before the majority-voting procedure described in Sec. 2.3.1).
Appendix B Hyperparameter/Ablation Studies
B.1 Object Tree-likeness Experiments
B.1.1 Synthetic Data Experiments
Tables 4 and 5 show results on the synthetic tree-like dataset using a wide range of CPR and DoGD hyperparameters ($R$ and $a,b$, respectively).
ViT-H ViT-B
$R$ $\tau$ $\rho$ $\tau$ $\rho$
$1$ $-0.82$ $-0.96$ $-0.79$ $-0.94$
$3$ $-0.80$ $-0.95$ $-0.76$ $-0.94$
$5$ $-0.77$ $-0.93$ $-0.75$ $-0.93$
$7$ $-0.74$ $-0.92$ $-0.73$ $-0.91$
$9$ $-0.72$ $-0.91$ $-0.70$ $-0.90$
$11$ $-0.71$ $-0.90$ $-0.70$ $-0.89$
Table 4: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by CPR with different hyperparameter values $R$ on the synthetic dataset, to supplement Table 1.
ViT-H ViT-B
$a$ $b$ $\tau$ $\rho$ $\tau$ $\rho$
$63$ $3$ $0.42$ $0.60$ $0.48$ $0.66$
$63$ $7$ $0.45$ $0.63$ $0.49$ $0.68$
$63$ $15$ $0.44$ $0.64$ $0.48$ $0.68$
$63$ $31$ $0.45$ $0.66$ $0.49$ $0.71$
$127$ $3$ $0.61$ $0.80$ $0.66$ $0.84$
$127$ $7$ $0.63$ $0.81$ $0.67$ $0.84$
$127$ $15$ $0.62$ $0.81$ $0.66$ $0.84$
$127$ $31$ $0.60$ $0.80$ $0.64$ $0.83$
$255$ $3$ $0.66$ $0.84$ $0.68$ $0.85$
$255$ $7$ $0.66$ $0.84$ $0.68$ $0.86$
$255$ $15$ $0.65$ $0.84$ $0.67$ $0.85$
$255$ $31$ $0.59$ $0.79$ $0.60$ $0.80$
Table 5: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by DoGD with different hyperparameter values $a,b$ on the synthetic dataset, to supplement Table 1.
B.1.2 Real Data Experiments
Tables 6, 7, 8, and 9 show results on DIS and iShape using a wide range of CPR and DoGD hyperparameters ($R$ and $a,b$, respectively).
ViT-H ViT-B
$R$ $\tau$ $\rho$ $\tau$ $\rho$
$1$ $-0.56$ $-0.73$ $-0.61$ $-0.79$
$3$ $-0.55$ $-0.72$ $-0.61$ $-0.79$
$5$ $-0.59$ $-0.76$ $-0.63$ $-0.81$
$7$ $-0.53$ $-0.7$ $-0.59$ $-0.77$
$9$ $-0.53$ $-0.71$ $-0.57$ $-0.75$
$11$ $-0.51$ $-0.69$ $-0.56$ $-0.74$
Table 6: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by CPR with different hyperparameter values $R$ on DIS, to supplement Table 2.
antenna branch fence hanger log wire
$R$ SAM Enc. $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
1 ViT-H $-0.18$ $-0.25$ $-0.63$ $-0.83$ $-0.68$ $-0.88$ $-0.30$ $-0.43$ $-0.60$ $-0.78$ $-0.08$ $-0.12$
ViT-B $-0.19$ $-0.29$ $-0.64$ $-0.83$ $-0.68$ $-0.88$ $-0.47$ $-0.66$ $-0.61$ $-0.79$ $-0.12$ $-0.18$
3 ViT-H $-0.20$ $-0.30$ $-0.63$ $-0.83$ $-0.65$ $-0.85$ $-0.28$ $-0.40$ $-0.61$ $-0.79$ $-0.07$ $-0.10$
ViT-B $-0.17$ $-0.26$ $-0.64$ $-0.83$ $-0.65$ $-0.85$ $-0.48$ $-0.66$ $-0.61$ $-0.79$ $-0.13$ $-0.19$
5 ViT-H $-0.13$ $-0.21$ $-0.61$ $-0.81$ $-0.61$ $-0.81$ $-0.25$ $-0.37$ $-0.60$ $-0.79$ $-0.06$ $-0.10$
ViT-B $-0.18$ $-0.27$ $-0.63$ $-0.82$ $-0.63$ $-0.83$ $-0.44$ $-0.62$ $-0.62$ $-0.80$ $-0.12$ $-0.17$
7 ViT-H $-0.16$ $-0.23$ $-0.57$ $-0.77$ $-0.54$ $-0.74$ $-0.20$ $-0.29$ $-0.60$ $-0.79$ $-0.06$ $-0.10$
ViT-B $-0.15$ $-0.24$ $-0.59$ $-0.77$ $-0.55$ $-0.75$ $-0.39$ $-0.55$ $-0.62$ $-0.81$ $-0.10$ $-0.15$
9 ViT-H $-0.16$ $-0.23$ $-0.52$ $-0.72$ $-0.47$ $-0.65$ $-0.13$ $-0.19$ $-0.59$ $-0.78$ $-0.05$ $-0.08$
ViT-B $-0.16$ $-0.22$ $-0.57$ $-0.76$ $-0.49$ $-0.67$ $-0.29$ $-0.42$ $-0.62$ $-0.81$ $-0.10$ $-0.15$
11 ViT-H $-0.08$ $-0.13$ $-0.49$ $-0.67$ $-0.51$ $-0.70$ $-0.02$ $-0.03$ $-0.60$ $-0.78$ $-0.03$ $-0.04$
ViT-B $-0.12$ $-0.16$ $-0.54$ $-0.73$ $-0.55$ $-0.75$ $-0.21$ $-0.28$ $-0.64$ $-0.82$ $-0.07$ $-0.11$
Table 7: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by CPR with different hyperparameter values $R$ on iShape, to supplement Table 2.
ViT-H ViT-B
$a$ $b$ $\tau$ $\rho$ $\tau$ $\rho$
$63$ $3$ $0.63$ $0.83$ $0.57$ $0.78$
$63$ $7$ $0.59$ $0.79$ $0.57$ $0.77$
$63$ $15$ $0.54$ $0.72$ $0.50$ $0.67$
$63$ $31$ $0.45$ $0.62$ $0.39$ $0.56$
$127$ $3$ $0.58$ $0.77$ $0.45$ $0.64$
$127$ $7$ $0.58$ $0.77$ $0.49$ $0.66$
$127$ $15$ $0.51$ $0.67$ $0.42$ $0.57$
$127$ $31$ $0.36$ $0.50$ $0.26$ $0.39$
$255$ $3$ $0.53$ $0.71$ $0.42$ $0.59$
$255$ $7$ $0.53$ $0.71$ $0.42$ $0.59$
$255$ $15$ $0.47$ $0.66$ $0.36$ $0.52$
$255$ $31$ $0.29$ $0.43$ $0.19$ $0.29$
Table 8: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by DoGD with different hyperparameter values $a,b$ on DIS, to supplement Table 2.
antenna branch fence hanger log wire
$a$ $b$ SAM Enc. $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
63 3 ViT-H $0.39$ $0.55$ $0.62$ $0.82$ $0.65$ $0.85$ $-0.15$ $-0.23$ $0.23$ $0.35$ $0.35$ $0.50$
ViT-B $0.25$ $0.35$ $0.62$ $0.81$ $0.60$ $0.81$ $-0.07$ $-0.11$ $0.17$ $0.25$ $0.36$ $0.51$
63 7 ViT-H $0.28$ $0.38$ $0.61$ $0.81$ $0.62$ $0.83$ $-0.13$ $-0.21$ $0.22$ $0.33$ $0.34$ $0.49$
ViT-B $0.13$ $0.18$ $0.61$ $0.81$ $0.60$ $0.80$ $-0.07$ $-0.11$ $0.14$ $0.22$ $0.33$ $0.48$
63 15 ViT-H $0.10$ $0.12$ $0.59$ $0.78$ $0.59$ $0.80$ $-0.05$ $-0.07$ $0.19$ $0.30$ $0.29$ $0.43$
ViT-B $0.04$ $0.03$ $0.59$ $0.77$ $0.56$ $0.77$ $-0.01$ $-0.01$ $0.09$ $0.16$ $0.25$ $0.36$
63 31 ViT-H $0.06$ $0.04$ $0.55$ $0.74$ $0.46$ $0.65$ $0.07$ $0.10$ $0.14$ $0.26$ $0.09$ $0.13$
ViT-B $-0.11$ $-0.17$ $0.56$ $0.74$ $0.45$ $0.64$ $0.07$ $0.11$ $0.02$ $0.08$ $0.02$ $0.03$
127 3 ViT-H $0.23$ $0.30$ $0.52$ $0.70$ $0.46$ $0.65$ $0.52$ $0.72$ $0.30$ $0.46$ $0.08$ $0.11$
ViT-B $0.12$ $0.14$ $0.51$ $0.69$ $0.45$ $0.64$ $0.50$ $0.70$ $0.13$ $0.22$ $0.06$ $0.08$
127 7 ViT-H $0.23$ $0.26$ $0.48$ $0.65$ $0.41$ $0.59$ $0.51$ $0.70$ $0.29$ $0.46$ $-0.01$ $-0.01$
ViT-B $0.10$ $0.09$ $0.51$ $0.68$ $0.41$ $0.59$ $0.49$ $0.68$ $0.13$ $0.22$ $-0.03$ $-0.05$
$127$ $15$ ViT-H $0.14$ $0.17$ $0.43$ $0.59$ $0.38$ $0.55$ $0.50$ $0.69$ $0.28$ $0.43$ $-0.09$ $-0.12$
ViT-B $0.07$ $0.03$ $0.45$ $0.62$ $0.36$ $0.52$ $0.48$ $0.67$ $0.12$ $0.20$ $-0.14$ $-0.20$
$127$ $31$ ViT-H $0.05$ $0.05$ $0.36$ $0.50$ $0.08$ $0.12$ $0.48$ $0.67$ $0.30$ $0.45$ $-0.15$ $-0.22$
ViT-B $0.04$ $0.00$ $0.38$ $0.52$ $0.12$ $0.18$ $0.46$ $0.65$ $0.12$ $0.20$ $-0.21$ $-0.31$
$255$ $3$ ViT-H $0.04$ $0.06$ $0.33$ $0.46$ $-0.20$ $-0.30$ $0.44$ $0.61$ $0.52$ $0.72$ $-0.17$ $-0.25$
ViT-B $0.21$ $0.31$ $0.30$ $0.42$ $-0.13$ $-0.20$ $0.42$ $0.59$ $0.42$ $0.60$ $-0.23$ $-0.34$
$255$ $7$ ViT-H $0.02$ $0.02$ $0.29$ $0.41$ $-0.25$ $-0.35$ $0.43$ $0.60$ $0.51$ $0.71$ $-0.20$ $-0.29$
ViT-B $0.19$ $0.27$ $0.26$ $0.36$ $-0.18$ $-0.26$ $0.42$ $0.58$ $0.41$ $0.59$ $-0.25$ $-0.36$
$255$ $15$ ViT-H $-0.04$ $-0.06$ $0.24$ $0.34$ $-0.31$ $-0.45$ $0.43$ $0.61$ $0.50$ $0.70$ $-0.20$ $-0.30$
ViT-B $0.21$ $0.29$ $0.20$ $0.28$ $-0.26$ $-0.39$ $0.40$ $0.56$ $0.40$ $0.58$ $-0.28$ $-0.40$
$255$ $31$ ViT-H $-0.11$ $-0.14$ $0.11$ $0.16$ $-0.47$ $-0.65$ $0.43$ $0.59$ $0.46$ $0.66$ $-0.23$ $-0.33$
ViT-B $0.13$ $0.20$ $0.07$ $0.09$ $-0.42$ $-0.60$ $0.39$ $0.55$ $0.38$ $0.55$ $-0.30$ $-0.44$
Table 9: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by DoGD with different hyperparameter values $R$ on iShape, to supplement Table 2.
B.2 Object Textural Separability Experiments
Table 10 show the textural separability experiment results on real data (iShape and Plittersdorf) using various hyperparameters and models for the classifier $g$.
iShape
antenna branch fence hanger log wire Plittersdorf
Weak Classifier
Model
SAM Enc. $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
Logistic, $C=2$ ViT-H $0.44$ $0.59$ $0.50$ $0.68$ $0.65$ $0.85$ $0.63$ $0.83$ $0.33$ $0.48$ $0.49$ $0.67$ $0.26$ $0.38$
ViT-B $0.46$ $0.61$ $0.56$ $0.75$ $0.67$ $0.86$ $0.65$ $0.84$ $0.36$ $0.52$ $0.55$ $0.73$ $0.36$ $0.50$
Logistic, $C=1$ ViT-H $0.42$ $0.58$ $0.49$ $0.67$ $0.64$ $0.84$ $0.62$ $0.81$ $0.31$ $0.46$ $0.49$ $0.67$ $0.34$ $0.49$
ViT-B $0.44$ $0.60$ $0.55$ $0.75$ $0.66$ $0.85$ $0.63$ $0.82$ $0.35$ $0.50$ $0.55$ $0.72$ $0.39$ $0.55$
Logistic, $C=0.5$ ViT-H $0.39$ $0.53$ $0.49$ $0.67$ $0.65$ $0.85$ $0.60$ $0.80$ $0.30$ $0.43$ $0.48$ $0.65$ $0.33$ $0.47$
ViT-B $0.48$ $0.65$ $0.57$ $0.76$ $0.66$ $0.86$ $0.62$ $0.81$ $0.33$ $0.48$ $0.52$ $0.69$ $0.41$ $0.56$
Random Forest, $n_{\mathrm{estimators}}=2$ ViT-H $0.48$ $0.66$ $0.45$ $0.63$ $0.46$ $0.64$ $0.45$ $0.64$ $0.19$ $0.27$ $0.25$ $0.36$ $0.35$ $0.50$
ViT-B $0.50$ $0.67$ $0.49$ $0.68$ $0.47$ $0.65$ $0.50$ $0.69$ $0.24$ $0.34$ $0.31$ $0.45$ $0.41$ $0.56$
Random Forest, $n_{\mathrm{estimators}}=4$ ViT-H $0.51$ $0.69$ $0.42$ $0.60$ $0.54$ $0.73$ $0.48$ $0.67$ $0.21$ $0.30$ $0.30$ $0.43$ $0.22$ $0.31$
ViT-B $0.51$ $0.70$ $0.48$ $0.66$ $0.53$ $0.72$ $0.51$ $0.70$ $0.22$ $0.32$ $0.37$ $0.52$ $0.32$ $0.45$
Random Forest, $n_{\mathrm{estimators}}=8$ ViT-H $0.48$ $0.65$ $0.46$ $0.63$ $0.56$ $0.75$ $0.50$ $0.69$ $0.19$ $0.28$ $0.35$ $0.50$ $0.22$ $0.31$
ViT-B $0.49$ $0.67$ $0.49$ $0.68$ $0.55$ $0.74$ $0.54$ $0.73$ $0.21$ $0.31$ $0.40$ $0.57$ $0.25$ $0.35$
Random Forest, $n_{\mathrm{estimators}}=16$ ViT-H $0.45$ $0.62$ $0.48$ $0.66$ $0.57$ $0.77$ $0.52$ $0.72$ $0.20$ $0.29$ $0.38$ $0.53$ $0.08$ $0.12$
ViT-B $0.58$ $0.77$ $0.52$ $0.70$ $0.54$ $0.74$ $0.54$ $0.74$ $0.23$ $0.33$ $0.41$ $0.57$ $0.15$ $0.23$
Table 10: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object textural separability (Algorithm 3), on real datasets (iShape and Plittersdorf), using different weak classifier models $g$ and hyperparameters to measure textural separability (to supplement Table 3).
Appendix C Additional Experimental Details
C.1 SAM Prompting Strategies
In addition to the SAM prompting details presented in Sec. 1, all real and style-transferred (NST) images are prompted with (1) a tight bounding-box and (2) several positive and/or negative prompts randomly sampled from either the foreground or background of the object mask, respectively, with respective counts $n_{pos}$ and $n_{neg}$. For iShape, we use $n_{pos}=n_{neg}=5$; for DIS, we use $n_{pos}=5$ and $n_{neg}=10$ due to the complexity of the objects. For Plittersdorf and NST images, we simply use $n_{pos}=0$ and $n_{neg}=2$, due to the objects typically possessing simple shapes.
C.2 Neural Style Transfer Experiments
Models.
For the NST experiments (Sec. 3.2.1), the inpainting model used is Runway’s stable diffusion inpainting pipeline, fp16 variant. The NST model itself is an implementation of [15] based on https://github.com/pytorch/tutorials/blob/main/advanced_source/neural_style_tutorial.py, using typical settings of a standard VGG19 CNN with content layer of conv_4, style layers of conv_i for i = 1, 2, 3, 4, 5, and ImageNet normalization.
Style transfer intensity.
As mentioned in Sec. 3.2.1, we perform style transfer experiments for a monotonically-increasing range of eight degrees of style transfer intensity. This is created by defining content and style weights $\lambda_{c}$ and $\lambda_{s}$, respectively for the NST algorithm according to
$\lambda_{c}=\frac{1}{1+\alpha}\quad\text{and}\quad\lambda_{s}=\frac{\alpha}{1+\alpha},$ (6)
where $\alpha$ ranges linearly from $\alpha=1$ to $\alpha=4,000$ with eight equally-spaced values, representing the degree of style transfer intensity.
Object/image altering procedure.
The procedure that creates the “altered” version of objects via non-affine transformations is detailed as follows.
  1. 1.
    Apply a non-affine transformation created using the albumentations Python library [5] to both the image and the object/mask, defined shortly.
  2. 2.
    Clean up unexpected obsolete regions.
  3. 3.
    Properly align the location of the distorted object.
  4. 4.
    Remove small isolated regions of the object.
In Python code, this is implemented as follows, given an input image NumPy array raw_img_arr and corresponding binary object mask raw_msk_arr:
import torch
import numpy as np
from PIL import Image
import albumentations as A
import torchvision.transforms as transforms
from skimage.morphology import dilation,label, \
disk, remove_small_holes, remove_small_objects, \
opening, closing
shape_transformation = A.Compose([
A.ElasticTransform(
alpha=2500, sigma=24, alpha_affine=0,
interpolation=1, border_mode=0, value=0,
mask_value=0, always_apply=True,
approximate=False, same_dxdy=False, p=1.0),
A.GridDistortion(
num_steps=20, distort_limit=.99,
interpolation=1, border_mode=0, value=0,
mask_value=0, normalized=False,
always_apply=False, p=1.0)]
imsize = 512
loader = transforms.Compose([
transforms.Resize(imsize), # scale imported image
transforms.CenterCrop(imsize),
transforms.ToTensor()])
unloader = transforms.ToPILImage()
def shift_array(array, shift_spec):
assert len(shift_spec) == 2 and \
isinstance(array,np.ndarray)
shift_corrected_arr = array.copy()
if shift_spec[0] != 0:
shift = shift_spec[0]
shift_corrected_arr = np.concatenate(
[shift_corrected_arr[shift:],
shift_corrected_arr[:shift]],
axis=0)
if shift_spec[1] != 0:
shift = shift_spec[1]
shift_corrected_arr = np.concatenate(
[shift_corrected_arr[:,shift:],
shift_corrected_arr[:,:shift]],
axis=1)
return shift_corrected_arr
def center_correction(raw_msk, distorted_msk):
assert raw_msk.ndim == 4 and \
distorted_msk.ndim == 3
pos_idx = torch.argwhere(raw_msk[0,0])
_raw_msk_center = (pos_idx.min(dim=0).values + \
pos_idx.max(dim=0).values)//2
pos_idx = torch.argwhere(distorted_msk[0])
_distorted_msk_center = (pos_idx.min(dim=0).values + \
pos_idx.max(dim=0).values)//2
position_correction = _distorted_msk_center -_raw_msk_center
corrected_distorted_msk = distorted_msk.squeeze()
print(corrected_distorted_msk.shape)
assert position_correction.ndim == 1 and len(position_correction) == 2
if position_correction[0] !=0:
shift = position_correction[0]
corrected_distorted_msk = torch.cat(
[corrected_distorted_msk[shift:],
corrected_distorted_msk[:shift]],
dim=0)
if position_correction[1] !=0:
shift = position_correction[1]
corrected_distorted_msk = torch.cat(
[corrected_distorted_msk[:,shift:],
corrected_distorted_msk[:,:shift]],
dim=1)
return corrected_distorted_msk
def get_distorted_ImageAndMask(trans, img, *msks):
# format check
assert all([msk.ndim == 2 for msk in msks])
# apply transformation to both image and msk
transformed = trans(image=img, masks = msks)
distorted_img_shifted = transformed[’image’]
distorted_msks_shifted = transformed[’masks’]
#clean up the unexpected obsolete regions
## get area excluding border
positive_region_map = label(distorted_msks_shifted[1])
label_ids, label_cnts = np.unique(positive_region_map,return_counts=True)
max_id = label_ids[1:][np.argmax(label_cnts[1:])]
eligible_region_map = np.where(positive_region_map==max_id, 1, 0)
distorted_msk_shifted = distorted_msks_shifted[0]*eligible_region_map
distorted_msk_shifted = remove_small_objects(remove_small_holes(
closing(opening(distorted_msk_shifted>0, disk(2)),disk(2))
#find center shift as a results of transformation
pos_idx = np.argwhere(msks[0]).squeeze()
_raw_msk_center = (pos_idx.min(0)+pos_idx.max(0))//2
pos_idx = np.argwhere(distorted_msk_shifted).squeeze()
_distorted_msk_center = (pos_idx.min(0)+pos_idx.max(0))//2
shift_spec = _distorted_msk_center -_raw_msk_center
#shift the arrays so that the obj location aligns
distorted_img = shift_array(distorted_img_shifted,shift_spec)
distorted_msk = shift_array(distorted_msk_shifted,shift_spec)
# remove relatively small regions
distorted_msk_labeled = label(distorted_msk)
unique_ids, id_cnts = np.unique(
distorted_msk_labeled, return_counts = True)
area_distribution = id_cnts[1:].astype(’float’)/id_cnts[1:].sum()
for area_id, area_rate in zip(unique_ids[1:], area_distribution):
if area_rate < 0.125:
distorted_msk_labeled[distorted_msk_labeled==area_id] = 0
distorted_msk_arr = distorted_msk_labeled>0
return distorted_img, distorted_msk_arr
distorted_img_arr, distorted_msk_arr = get_distorted_ImageAndMask(
shape_transformation,raw_img_arr,raw_msk_arr,raw_msk_arr.copy())
Appendix D Additional Results
D.1 Correlation Between Metrics
In Fig. D.1, we show the relationship between our proposed tree-likeness metrics CPR and DoGD on all three datasets (with default hyperparameters and ViT-H SAM), quantified by correlations shown in Table 11.
Figure 13: Relationship between tree-likeness metrics CPR and DoGD. From left to right, on synthetic data (Fig. 4), DIS, and iShape (Fig. 5).
iShape
Synth. DIS antenna branch fence hanger log wire all Avg.
Pearson $r$ $-0.94$ $-0.67$ $-0.92$ $-0.93$ $-0.82$ $-0.81$ $-0.90$ $-0.66$ $-0.79$ $-0.83$
Spearman $\rho$ $-0.82$ $-0.70$ $-0.91$ $-0.93$ $-0.86$ $-0.76$ $-0.94$ $-0.79$ $-0.76$ $-0.83$
Kendall $\tau$ $-0.62$ $-0.51$ $-0.75$ $-0.78$ $-0.68$ $-0.58$ $-0.79$ $-0.61$ $-0.57$ $-0.65$
Table 11: Linear (Pearson $r$) and nonlinear (Spearman $\rho$, Kendall $\tau$) correlations between CPR and DoGD on all evaluated datasets.

### v2

@@
-mO Haoyu[#1]
+Quantifying the Limits of Segmentation Foundation Models: Modeling Challenges in Segmenting Tree-Like and Low-Contrast Objects
 Abstract
-Segment Anything Model (SAM) has shown impressive performance in interactive and zero-shot segmentation across diverse domains, suggesting that they have learned a general concept of “objects” from their large-scale training.
-However, we observed that SAM struggles with certain types of objects, particularly those featuring dense, tree-like structures and low textural contrast from their surroundings.
-These failure modes are critical for understanding its limitations in real-world use.
-In order to systematically examine this issue, we propose metrics to quantify two key object characteristics: tree-likeness and textural separability.
-Through extensive controlled synthetic experiments and testing on real datasets, we demonstrate that SAM’s performance is noticeably correlated with these factors.
-We link these behaviors under the concept of “textural confusion”, where SAM misinterprets local structure as global texture, leading to over-segmentation, or struggles to differentiate objects from similarly textured backgrounds.
-These findings offer the first quantitative framework to model SAM’s challenges, providing valuable insights into its limitations and guiding future improvements for vision foundation models.
-Introduction
-Segment Anything Model (SAM) [27], the most widely-used foundation model for image segmentation, has demonstrated promising zero-shot and fine-tuned segmentation ability for objects not seen in training, within a wide range of domains since its release, such as biomedical images [30, 22, 10], remote sensing [6, 35], etc.
-However, certain failure modes for SAM have been found empirically, where SAM produces surprisingly underwhelming performance on datasets with specific uncommon objects or atypical contexts, such as those with dense, branching structures (e.g., retinal blood vessels) [36, 11], concealed/low-contrast objects [7, 39, 24], small or irregular objects, and others [24] [^1] We focus on “unintentional” failure modes of SAM, as opposed to intentional adversarial attacks/examples [45, 47]..
-A solution to this problem could simply be fine-tuning the model on such challenging examples using various strategies [29, 32, 17], but this would still lack an understanding of the underlying causes of the issues, and so may not result in the best solution.
-As such, it is important to develop a better theory for why these failures happen, in order to better understand (and potentially mitigate) the limitations of SAM for new applications.
+Image segmentation foundation models (SFMs) like Segment Anything Model (SAM) have achieved impressive zero-shot and interactive segmentation across diverse domains.
+However, they struggle to segment objects with certain structures, particularly those with dense, tree-like morphology and low textural contrast from their surroundings.
+These failure modes are crucial for understanding the limitations of SFMs in real-world applications.
+To systematically study this issue, we introduce interpretable metrics quantifying object tree-likeness and textural separability.
+On carefully controlled synthetic experiments and real-world datasets, we show that SFM performance (e.g., SAM, SAM 2, HQ-SAM) noticeably correlates with these factors.
+We link these failures to “textural confusion”, where models misinterpret local structure as global texture, causing over-segmentation or difficulty distinguishing objects from similar backgrounds.
+Notably, targeted fine-tuning fails to resolve this issue, indicating a fundamental limitation.
+Our study provides the first quantitative framework for modeling the behavior of SFMs on challenging structures, offering interpretable insights into their segmentation capabilities.
+1 Introduction
+Segment Anything Model (SAM) [26], the most widely-used segmentation foundation model (SFM), has demonstrated promising zero-shot and fine-tuned segmentation ability across diverse domains, including biomedical imaging [31, 21, 10], remote sensing [6, 37], and others.
+However, certain failure modes for SAM have been found empirically, where SAM produces underwhelming performance on datasets with certain uncommon objects or atypical contexts, such as dense, branching structures (e.g., retinal blood vessels) [38, 11], concealed/low-contrast objects [7, 41, 23, 44], small or irregular objects, and others [23].
+We will show that newer SFMs, including SAM 2 [36] and HQ-SAM [24], are also susceptible to these failure modes, despite some stated improvements over these limitations for SAM—HQ-SAM on thin, branching structures and SAM 2 on low-contrast objects.
+Fine-tuning or adapting these models for such challenging examples using various strategies [30, 33, 16, 24] (or utilizing specialized techniques, e.g., matting algorithms [28]) could mitigate these issues, yet we will show that these strategies do not fully resolve the problems, nor explain the root causes.
+This underscores the need for a theoretical understanding of why SFMs struggle with such cases, paving the way for better model design and adaptation for new applications.
 Figure 1: SAM’s segmentation performance tends to drop noticeably when the object has high tree-likeness (left, on DIS) or low textural separability (right, on iShape)–even with heavy prompting–which we investigate in this work.
 From left to right: input images, object annotations, and their SAM segmentations.
-In this paper, we propose that several of these failure modes can be explained as SAM misinterpreting textural cues for how it disambiguates an object from its surroundings: it typically confuses objects with irregular, dense tree-like local structure as large-scale textures (resulting in over-segmentation), and similarly is generally challenged by objects that have similar textures to their surroundings.
-We propose metrics developed from first principles for quantifying the characteristics of (1) object tree-likeness and (2) textual separability.
-Objects tree-likeness (Sec.
-2.2) is based on concepts such as spatial contour frequency and the difference in typical object variability between global and local scales.
+Here, we propose that several of these failure modes in SFMs stem from misinterpreting textural cues when distinguishing objects from their surroundings.
+Specifically, SFMs often confuse dense, tree-like structures as large-scale textures, leading to over-segmentation, and struggle with objects that have similar textures to their surroundings.
+To quantify these challenges, we introduce interpretable metrics for object tree-likeness and textural separability.
+Tree-likeness (Sec.
+4.2) captures structural complexity using spatial contour frequency and the difference in object variability between global and local scales.
 Textural separability (Sec.
-3.1) is based on the detectability of differences in textural features of an object compared to its surroundings.
-We designed these metrics to be simple and interpretable, such that they can be applied to any dataset or object type.
-We also present them in quickly-computable PyTorch-like algorithmic form that can be applied to any masks.
-Next, we begin to probe the effect that these characteristics have on SAM’s segmentation ability through experiments on synthetic data which are designed to cover a range of object tree-likeness (Sec.
-2.3.1) and textural separability (Sec.
-3.2.1).
-These experiments are carefully controlled to disentangle the effects of object shape and texture on performance, as well as mitigate any other potential confounding factors.
-We then continue with similar experiments on real datasets (Secs.
-2.3.2 and 3.2.2) that represent a wide range of objects to segment.
-In all experiments, we find both the exact measured tree-likeness and the textural separability of objects are noticeably correlated with SAM’s performance (IoU) in segmenting it, over a wide range of object and image types.
-Our findings are the first to quantitatively model and experimentally verify how the performance of segmentation foundation models is affected by certain measurable object characteristics, providing an explanation for prior findings of SAM’s unexpected failures on new datasets.
+5.1) measures how distinguishable an object’s textural features are from those of its background.
+These metrics are designed for broad applicability across datasets and object types and are implemented in a PyTorch-compatible form for efficiency.
+We then investigate how tree-likeness and textural separability affect SFM performance through controlled synthetic experiments (Secs.
+4.3.1, 5.2.1) and real-world datasets (Secs.
+4.3.2, 5.2.2).
+These experiments systematically isolate the influence of object shape and texture while minimizing confounding factors.
+Across all datasets, we find that SFM performance (IoU) noticeably correlates with both tree-likeness and textural separability across diverse object and image types.
+Notably, HQ-SAM, designed for intricate structures, still exhibits these limitations, highlighting the persistence of these challenges.
+Furthermore, our fine-tuning experiments (Sec.
+6) confirm that these failure modes remain robust to fine-tuning, even when explicitly trained on the most difficult cases.
 Our overall contributions are as follows:
   1. 1.
-    Based on first principles, we propose quantitative metrics for object tree-likeness of Contour Pixel Rate (CPR) and Difference of Gini Impurity Deviation (DoGD), as well as a metric for object textural separability.
+    We introduce quantitative metrics for object tree-likeness—Contour Pixel Rate (CPR) and Difference of Gini Impurity Deviation (DoGD)—and an object textural separability metric based on neural feature contrast.
   2. 2.
-    We present carefully-designed experiments for the segmentation of synthetic images which cover a range of tree-likeness and textural separability, where important factors such as object shape and texture are controlled.
+    Through carefully controlled synthetic experiments and real datasets, we demonstrate that the segmentation performance of SAM, SAM 2, and HQ-SAM negatively correlates with tree-likeness and positively correlates with textural separability.
   3. 3.
-    We show that SAM’s segmentation performance tends to negatively correlate with an object’s tree-likeness, and positively correlate with it’s textural separability, on a variety of experiments on both synthetic and real images.
-We hope that this inspires future work in understanding the limitations of vision foundation models.
+    We show that these segmentation challenges persist even after fine-tuning, suggesting that SFMs struggle with these object types regardless of adaptation.
+Our findings provide the first quantitative model linking SFM performance to measurable object characteristics.
+Despite the complexity of SFMs and their evaluation datasets, we show that their underperformance on challenging cases can be modeled using our straightforward, interpretable object descriptors.
 We release easy-to-use code for our proposed metrics at https://github.com/mazurowski-lab/SAM-TexturalConfusion-Metrics.
-Related Works
+2 Related Work
 In addition to the various “failure modes” and biases of segmentation foundation models discussed in the introduction which motivated this work, various works have also studied the same for general vision models.
@@
 For example, [27] found that when a CNN is presented with images of objects with only their shape visible, not texture (silhouettes), performance was greatly worsened.
-It has also been found that the shape-learning behavior of neural networks can differ noticeably depending on the particular training dataset and objective [21, 20], motivating us to study the specific behavior of SAM (given its’ unique training dataset and pipeline) in detail.
+It has also been found that the shape-learning behavior of neural networks can differ noticeably depending on the particular training dataset and objective [20, 19], motivating us to study the specific behavior of SAM (given its unique training set) and other similar SFMs in detail.
 Interestingly, there has been some evidence that vision transformers (like SAM’s image encoder) may instead be biased towards shape, rather than texture for classification tasks [42].
-Here, we study segmentation rather than classification, where it is unclear if those results extend, as our results generally point to SAM focusing on texture over shape (to the point of even confusing complex, locally-varying shape itself as texture).
-1 Methods: SAM Usage and Prompting
-We experiment with both the ViT-H (default) and ViT-B standard pretrained SAM models [27].
-As is default for SAM, all input images are resized to a resolution of $1024\times 1024$, and normalized to $[0,255]$.
-All images that possess instance segmentations for multiple objects (e.g., iShape and Plittersdorf) will be evaluated by SAM segmenting each of the images’ objects one at a time (guided by clear prompting), with all other mask pixels set as background.
-In all experiments, we use the oracle-chosen mask out of SAM’s three outputted predictions.
+Here, we study segmentation rather than classification, where it is unclear if those results extend, as our results generally point to SAM focusing on texture over shape (to the point of even confusing complex, locally varying shape itself as texture), which aligns with the preliminary findings of [48].
+3 Methods: SFM Usage and Prompting
+We experiment with the ViT-H (default, if unstated) and ViT-B standard pre-trained SAM models [26], the ViT-L and ViT-B+ variants of SAM 2 [36], and the ViT-H and ViT-B variants of HQ-SAM [24], a SAM adaptation designed particularly for thin structures.
+As is the default for SAM, all input images are resized to a resolution of $1024\times 1024$, and normalized to $[0,255]$.
+All images that possess instance segmentations for multiple objects will be evaluated by the given SFM segmenting each of the images’ objects one at a time (guided by clear prompting described in the next paragraph), with all other mask pixels set as background.
+In all experiments, we use the oracle-chosen mask out of the given SFM’s three outputted predictions.
 For all experiments on the synthetic images of Sec.
-2.3.1, we provide a tight bounding box about the object of interest as the prompt.
+4.3.1, we provide a tight bounding box about the object of interest as the prompt.
 For all experiments on real images as well as the style-transferred images of Sec.
-3.2.1, we provide the same bounding box as well as a certain number of positive and/or negative point prompts randomly sampled from the object foreground and background within the bounding box, respectively, depending on the dataset (full details in Appendix C.1).
-We use this relatively heavy prompting strategy in order to minimize the ambiguity of instructions provided to SAM, which will also help to minimize any differences of the oracle prediction from the other two.
+5.2.1, we provide the same bounding box as well as a certain number of positive and/or negative point prompts randomly sampled from the object foreground and background within the bounding box, respectively, depending on the dataset (full details in Appendix C.1).
+We use this relatively heavy prompting strategy to minimize the ambiguity of instructions provided to the given SFM, which will also help to minimize any differences in the oracle prediction from the other two.
 4 The Challenge of Tree-Like Structures
 4.1 Motivation
-We first wished to understand SAM’s challenges in segmenting certain objects with dense, tree-like structures, following our observation of a trend of SAM having difficulty with such objects including retinal blood vessels [34, 29] or satellite road images [14, 42] (see e.g., Fig.
+We first wished to understand the challenges for SFMs in segmenting certain objects with dense, tree-like structures, and then quantify the features driving these challenges, following our observation of a trend of SAM having difficulty with such objects, including retinal blood vessels [35, 30] or satellite road images [13, 45] (see e.g., Fig.
 2).
-Interestingly, while both of these object types have branching features, a characteristic which naively could relate to SAM’s failure, SAM’s performance on retinal vessel images is noticeably worse (avg.
+Interestingly, while both of these object types have branching features, a characteristic which naively could relate to the SFM’s “failure”, SAM’s performance on retinal vessel images is noticeably worse (avg.
 $\mathrm{IoU}\simeq 0.05$ [35]) compared to on satellite road structures (avg.
 $\mathrm{IoU}\simeq 0.2$ [13]), which was reproduced in our own experiments (Fig.
-4).
+4) [^1] This could potentially simply be due to retinal vessels being thinner than road structures, but our experiments do not support this (App.
+D.2)..
 Figure 2: Example retinal blood vessel (top) and satellite road (bottom) images and accompanying object segmentation masks.
-This motivated us to quantify the features of these types of objects that relate to SAM’s failure.
 We hypothesize that such objects become more challenging to segment when their branching structures are dense and irregularly-spaced, which we refer to as “tree-like”.
-We will show that SAM considers such structures as textures rather than shapes, resulting in significant over-segmentation.
-In the following section we propose how to characterize the tree-likeness of objects using two new quantitative metrics: CPR (Contour Pixel Rate) and DoGD (Difference of Gini Impurity Deviation).
+We will show that SFMs consider such structures as textures rather than shapes, resulting in significant over-segmentation.
+In the following section, we propose how to characterize the tree-likeness of objects using two new quantitative metrics: CPR (Contour Pixel Rate) and DoGD (Difference of Gini Impurity Deviation).
 4.2 Quantifying Tree-Like Structures
@@
 Intuitively, objects with significant tree-like or fractal-like structure will exhibit relatively large values of $\sigma^{\mathrm{Gini}}_{b}(m)$ due to high variability in pixel composition at small scales (frequently alternating between areas of mixed classes and areas of a single class), yet small $\sigma^{\mathrm{Gini}}_{a}(m)$ due to structures with high uniformity and/or repetitions at large scales, altogether increasing the DoGD.
-We perform all experiments with the hyperparameters for CPR and DoGD set to $R=5$, $a=127$, and $b=3$, which we found via grid search for the values which resulted in the Kendall’s $\tau$ between IoU and DoGD with the lowest $p$-value on the held-out DIS training set using ViT-H SAM.
-We show results using a wide range of other values for these hyperparameters in Appendix B.1, where we found our findings to be consistent for most other settings.
-Moreover, we note that CPR and DoGD are correlated with each other (Pearson $|r|=0.83$ on average; Appendix D.1), showing their consistency in how they quantify different aspects of tree-likeness.
+We perform all experiments with the hyperparameters for CPR and DoGD set to $R=5$, $a=127$, and $b=3$, which we found via grid search for the values that resulted in the Kendall’s $\tau$ between IoU and DoGD with the lowest $p$-value on the held-out DIS training set using ViT-H SAM.
+We show results using a wide range of other values for these hyperparameters in Appendix B.1, where our findings were consistent for most other settings.
+Moreover, we note that CPR and DoGD are correlated with each other (Pearson $|r|=0.83$ on average; Appendix D.1), showing their consistency in quantifying different aspects of tree-likeness.
 4.3 The Relationship between Object Tree-likeness and Segmentation Performance
 4.3.1 Experiments on Synthetic Data
-In this section, we will first carefully probe the effect of object tree-likeness on SAM’s segmentation performance by testing it on synthetic images which solely possess objects of varying tree-likeness, with different independently chosen, uniform foreground and background textures.
-These objects are contiguous components samples from retinal blood vessel and satellite road masks, with the full procedure of generating these images detailed in Appendix A.2.
-Example generated images, masks and SAM segmentation predictions for them are shown in Fig.
+In this section, we will first probe the effect of object tree-likeness on the performance of SFMs by testing them on synthetic images that solely possess objects of varying tree-likeness, with different independently chosen, uniform foreground and background textures.
+These objects are contiguous components sampled from retinal blood vessel and satellite road masks; the full procedure of generating these images is detailed in Appendix A.2.
+Example generated images, masks, and SAM predictions for them are shown in Fig.
 3.
-As shown (as well in Fig.
+As shown (also in Fig.
 4), the objects cover a wide range of tree-likeness as measured by these quantities.
-In order to mitigate any confounding on SAM performance due to an object’s textural contrast from its surroundings (which we study in Sec.
-3), for each mask $m_{c}$ generated by our procedure, we apply SAM to $K=7$ images created by applying $K$ different randomly-sampled pairs of textures to the object’s foreground ($m_{c}=1$) and background ($m_{c}=0$).
-We then obtain SAM’s final prediction for this object via pixel-wise majority voting over its predictions on these $K$ images.
+In order to mitigate any confounding effects on a given SFM’s performance due to an object’s textural contrast from its surroundings (which we study in Sec.
+5), for each mask $m_{c}$ generated by our procedure, we apply the SFM to $K=7$ images created by applying $K$ different randomly-sampled pairs of textures to the object’s foreground ($m_{c}=1$) and background ($m_{c}=0$).
+We then obtain the SFM’s final prediction for this object via pixel-wise majority voting over its predictions on these $K$ images.
 Figure 3: Left: Example synthetic tree-like images and object masks.
-Right: Trend of increasing tree-likeness (increasing CPR/decreasing DoGD) of these objects resulting in worse SAM segmentation predictions (to the right of each object mask).
+Right: Trend of increasing tree-likeness (increasing CPR/decreasing DoGD) of these objects (left in each pair) resulting in worse SAM segmentation predictions (right in each pair)
 In Fig.
-4 we show the relationship between an object’s tree-likeness (as measured by CPR or DoGD) with SAM’s performance (IoU) in segmenting the object, for all generated synthetic objects.
-We quantify the strength of this relation via rank/non-linear correlation, measured by Kendall’s tau ($\tau$) [25] and Spearman’s rho ($\rho$) [37], shown in Table 1333We evaluate $\tau$ in addition to $\rho$ due to it being more robust to outliers..
-Intriguingly, we see that object tree-likeness is quite predictive of SAM’s performance, with average absolute correlations of $|\tau|=0.76$ and $|\rho|=0.93$ for CPR and $|\tau|=0.64$ and $|\rho|=0.82$ for DoGD; i.e., more prevalent dense tree-like structures corresponding to worse performance.
+4 we show the relationship between an object’s tree-likeness (via CPR or DoGD) with SAM’s performance (IoU) in segmenting the object, for all generated synthetic objects.
+We quantify the strength of this relation via non-linear correlation, measured by Kendall’s tau ($\tau$) [25] and Spearman’s rho ($\rho$) [39], for all three SFMs, shown in Table 1.
+We see that object tree-likeness is quite predictive of an SFM’s performance, with average absolute correlations of $|\tau|=0.75$ and $|\rho|=0.92$ for CPR and $|\tau|=0.67$ and $|\rho|=0.84$ for DoGD; i.e., more prevalent dense tree-like structures corresponding to worse performance.
 CPR DoGD
-SAM Encoder $\tau$ $\rho$ $\tau$ $\rho$
-ViT-H -0.77 -0.93 0.61 0.80
-ViT-B -0.75 -0.93 0.66 0.84
-Table 1: Rank correlation between SAM segmentation IoU and object tree-likeness (CPR and DoGD), on the synthetic dataset.
-Figure 4: SAM segmentation IoU vs. object tree-likeness (CPR, left; DoGD, right) on the synthetic dataset.
+Model $\tau$ $\rho$ $\tau$ $\rho$
+SAM, ViT-H [26] -0.77 -0.93 0.61 0.80
+SAM, ViT-B [26] -0.75 -0.93 0.66 0.84
+SAM 2, ViT-L [36] -0.73 -0.91 0.67 0.85
+SAM 2, ViT-B+ [36] -0.71 -0.90 0.69 0.86
+HQ-SAM, ViT-H [24] -0.74 -0.92 0.70 0.87
+HQ-SAM, ViT-B [24] -0.81 -0.95 0.66 0.84
+Average -0.75 -0.92 0.67 0.84
+Table 1: Rank correlation between SFM IoU and object tree-likeness (CPR and DoGD), on the synthetic dataset.
+Figure 4: SAM IoU vs. object tree-likeness (CPR, left; DoGD, right) on the synthetic dataset.
 While this relationship is certainly strong, it is still on synthetic, controlled data.
@@
 We will now perform the same analysis on two real datasets which contain objects which cover a spectrum of tree-likeness, DIS5k and iShape.
-DIS5k [33] (or “DIS” for short) is a dataset that contains extremely detailed segmentation masks of objects with varying degrees of hollowness, and both regular and irregular tree-like structures.
+DIS5k [34] (or “DIS”) is a dataset containing extremely detailed annotations of objects with varying degrees of hollowness, and both regular and irregular tree-like structures.
 All DIS experiments will be reported on its validation set unless otherwise stated.
@@
 Fig.
-9).
+13).
 We analyze these classes individually to mitigate potential confounding/noise factors due to inter-class variations; we do not do this for DIS due to the larger number of classes which are much more fine-grained in their differences, additionally because there are few images per class for DIS.
-Results.
 In Fig.
-5 we show how SAM’s performance (IoU) on these images relates to the tree-likeness of the objects which it is segmenting, with accompanying quantitative correlation results shown in Table 2.
+5 we show how SAM’s performance (IoU) on these images relates to the tree-likeness of the objects which it is segmenting, alongside the other SFMs (SAM 2 and HQ-SAM) for comparison for DIS, with accompanying quantitative correlation results for all three SFMs shown in Table 2.
 In order to reduce the noise incurred by the large variety of segmented objects in the dataset and nuisance confounding factors in the images, we analyze results with aggregated objects: we cluster groups of $5$ objects/images with similar IoU and tree-likeness (CPR or DoGD) into single datapoints of the average value of these metrics, and similar for the IoU vs.
 textural separability experiments of Sec.
-3.
+5.
 iShape
 DIS antenna branch fence hanger log wire
-SAM Enc. $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
-CPR: ViT-H -0.59 -0.76 -0.13 -0.21 -0.61 -0.81 -0.61 -0.81 -0.25 -0.37 -0.60 -0.77 -0.06 -0.10
-ViT-B -0.63 -0.81 -0.18 -0.27 -0.63 -0.82 -0.63 -0.83 -0.44 -0.62 -0.62 -0.80 -0.12 -0.17
-DoGD: ViT-H 0.58 0.77 0.23 0.30 0.52 0.70 0.46 0.65 0.52 0.72 0.30 0.46 0.08 0.11
-ViT-B 0.45 0.64 0.12 0.14 0.51 0.69 0.45 0.64 0.50 0.70 0.13 0.22 0.06 0.08
-Table 2: Rank correlation between SAM prediction IoU and object tree-likeness (CPR, upper table) and DoGD, lower table), on DIS and iShape.
-Figure 5: SAM segmentation IoU vs. object tree-likeness (CPR, top; DoGD, bottom) on DIS (left) and iShape (right).
-We see that despite the many potential confounding factors in real data, there is still a clear correlation between object tree-likeness as measured by the proposed metrics and SAM segmentation performance.
-In particular, we see an average absolute correlations of $|\tau|=0.56$ and $|\rho|=0.74$ for CPR and $|\tau|=0.44$ and $|\rho|=0.62$ for DoGD, excluding the antenna and wire objects of iShape which had outlying correlations, likely because those two object classes cover only small ranges of tree-likeness as opposed to the other iShape classes (Fig.
+Model $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
+CPR: SAM, ViT-H [26] -0.59 -0.76 -0.13 -0.21 -0.61 -0.81 -0.61 -0.81 -0.25 -0.37 -0.60 -0.77 -0.06 -0.10
+SAM, ViT-B [26] -0.63 -0.81 -0.18 -0.27 -0.63 -0.82 -0.63 -0.83 -0.44 -0.62 -0.62 -0.80 -0.12 -0.17
+SAM 2, ViT-L [36] -0.41 -0.57 -0.30 -0.43 -0.64 -0.84 -0.61 -0.81 -0.48 -0.66 -0.67 -0.85 -0.14 -0.21
+SAM 2, ViT-B+ [36] -0.38 -0.52 -0.16 -0.23 -0.52 -0.71 -0.57 -0.78 -0.42 -0.59 -0.60 -0.77 -0.12 -0.18
+HQ-SAM, ViT-H [24] -0.52 -0.69 -0.23 -0.33 -0.77 -0.93 -0.79 -0.95 -0.51 -0.70 -0.69 -0.85 -0.22 -0.33
+HQ-SAM, ViT-B [24] -0.62 -0.81 -0.25 -0.34 -0.75 -0.92 -0.83 -0.96 -0.50 -0.69 -0.71 -0.87 -0.24 -0.35
+Average -0.53 -0.69 -0.21 -0.30 -0.65 -0.84 -0.67 -0.86 -0.43 -0.61 -0.65 -0.82 -0.15 -0.22
+DoGD: SAM, ViT-H [26] 0.58 0.77 0.23 0.30 0.52 0.70 0.46 0.65 0.52 0.72 0.30 0.46 0.08 0.11
+SAM, ViT-B [26] 0.45 0.64 0.12 0.14 0.51 0.69 0.45 0.64 0.50 0.70 0.13 0.22 0.06 0.08
+SAM 2, ViT-L [36] 0.44 0.60 0.42 0.59 0.48 0.65 0.49 0.68 -0.05 -0.08 0.56 0.75 -0.03 -0.05
+SAM 2, ViT-B+ [36] 0.26 0.39 0.40 0.56 0.29 0.40 0.30 0.44 -0.06 -0.10 0.42 0.58 -0.08 -0.12
+HQ-SAM, ViT-H [24] 0.48 0.66 0.53 0.72 0.58 0.72 0.48 0.67 -0.31 -0.45 0.57 0.75 -0.13 -0.19
+HQ-SAM, ViT-B [24] 0.50 0.69 0.67 0.85 0.59 0.73 0.57 0.77 -0.25 -0.34 0.53 0.71 -0.04 -0.06
+Average 0.45 0.63 0.40 0.53 0.50 0.65 0.46 0.64 0.06 0.08 0.42 0.58 -0.02 -0.04
+Table 2: Rank correlation between SFM IoU and object tree-likeness (CPR, upper table and DoGD, lower table), on DIS and iShape.
+Figure 5: Segmentation IoU vs. object tree-likeness (CPR, top; DoGD, bottom) on DIS (left) for all three SFMs and iShape (right) shown for SAM.
+We see that despite the many potential confounding factors in real data, there is still a clear correlation between object tree-likeness as measured by the proposed metrics and SFM segmentation performance.
+In particular, we see average absolute correlations of $|\tau|=0.59$ and $|\rho|=0.76$ for CPR and $|\tau|=0.38$ and $|\rho|=0.51$ for DoGD—excluding the antenna and wire objects of iShape which had outlying correlations likely because those two object classes cover only small ranges of tree-likeness as opposed to the other iShape classes (Fig.
 5), such that noise from other confounding factors obscures any dependence of performance on tree-likeness.
 5 The Challenge of Textural Separability
-In the previous section, we demonstrated that SAM struggles with segmenting non-conventional shapes, in particular, dense tree-like structures, which we hypothesize is due to the model confusing the dense structure as the texture of a non-treelike, more regular shape, rather than a shape itself (see e.g.
+In the previous section, we demonstrated that SFMs struggle with segmenting non-conventional shapes, in particular, dense tree-like structures, which we hypothesize is due to the model confusing the dense structure as the texture of a non-treelike, more regular shape, rather than a shape itself (see e.g.
 Fig.
 1 left).
-Similar to this behavior is that even for objects with simpler shapes, SAM can still be confused if the object’s texture is even somewhat similar to its surroundings, which we will study in this section.
+Similar to this behavior is that even for objects with simpler shapes, these SFMs can still be confused if the object’s texture is even somewhat similar to its surroundings, which we will study in this section.
 5.1 Measuring Textural Separability
 We will define the textural contrast or separability between some object mask and its surroundings by how easily their textures can be distinguished from one another.
-Motivated by findings that early layers of classification-pretrained CNNs primarily capture low-level features involving edges and textures [44], we will characterize an image’s textures using the first convolutional layer of a ResNet-18 [19] pretrained on ImageNet [9].
+Motivated by findings that early layers of classification-pretrained CNNs primarily capture low-level features involving edges and textures [47], we characterize an image’s textures using the first convolutional layer of a ResNet-18 [18] pretrained on ImageNet [9].
 Denoted by $f_{1}:\mathbb{R}^{C\times H\times W}\rightarrow\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$, this outputs a textural feature map corresponding to all $7\times 7$ windows in the image.
-We then measure the textural separability of an object according to if a simple classifier $g$ can be trained to discriminate between (a) the activations for the pixels of the object foreground and (b) the activations right outside of the object boundary [^4] This procedure is somewhat similar to the probing of hidden activation concepts [2, 26], although these works detected concepts (such as textures) at the image level, not at the single activation level as we do here..
+We then measure the textural separability of an object according to if a simple classifier $g$ can be trained to discriminate between (a) the activations for the pixels of the object foreground and (b) the activations right outside of the object boundary.
 We define this explicitly in Algorithm 3, where $\mathrm{Dilate}$ and $\mathrm{Disk}$ refer to scikit-image.morphology functions [43].
@@
 Algorithm 3 Textural separability of an object.
-0:  Image $x\in\mathbb{R}^{C\times H\times W}$ with mask $m\in\{0,1\}^{H\times W}$, first convolutional layer of pretrained CNN $f_{1}$, simple classifier $g$.
+0:  Image $x$ with binary mask $m$, first convolutional layer of pre-trained CNN $f_{1}$, simple classifier $g$.
 1:  $h=f_{1}(x)\in\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$ (Get textural features.)
@@
 4:  $m^{\prime}=m^{\prime}-m$ (Extract boundary only.)
-5:  $h_{\mathrm{obj}}:=h[:,m\neq 0]$
-6:  $h_{\mathrm{bdry}}:=h[:,m^{\prime}\neq 0]$
-7:  Train $g$ to classify between activations in $h_{\mathrm{obj}}$ vs. $h_{\mathrm{bdry}}$.
-8:  return  Training classification accuracy of $g$.
+5:  $h_{\mathrm{obj}}:=h[:,m\neq 0]$, $h_{\mathrm{bdry}}:=h[:,m^{\prime}\neq 0]$
+6:  Train $g$ to classify between activations in $h_{\mathrm{obj}}$ vs. $h_{\mathrm{bdry}}$.
+7:  return  Training classification accuracy of $g$.
 5.2 The Effect of Textural Separability on Segmentation Performance
 5.2.1 Style Transfer for Controlled Textural Contrast
-Similar to Sec.
-2.3, in studying the effects of textural separability on SAM’s performance, we also wish to perform experiments on well-controlled synthetic images in order to more carefully disentangle the effects of textural separability shape on performance from the effects due to object shape, before evaluating on real data (in the following section).
-Here, we will do so using neural style transfer [15] to precisely modify the textural contrast of real objects/images.
-This will involve (1) modifying objects’ shape without altering their background using an inpainting model, followed by (2) applying neural style transfer (NST) to the composite image to adjust the textural contrast of the object with the surrounding image.
-We begin with images sampled from the VOC2012 dataset [13] accompanied by objects (instance masks), according to certain criteria.
-Namely, we pick objects which take up between $5\%$ and $25\%$ of the entire image’s area (to have large enough objects while still maintaining sufficient background), sampling 484 objects total (with accompanying background) for study.
-For each object, we use a Stable Diffusion-based inpainting model (details in Appendix C.2) to remove the object from its corresponding image and fill in it’s mask area with the background.
-This creates a pair of a background image and a separate object (mask), allowing us to have careful control of modifying the object (and its mask accordingly) before placing it back into the inpainted background.
-For a given object, we then create three types of composite images with varying degrees of changes to the object’s shape (without modifying the background):
+To isolate the effects of textural separability from object shape, we first conduct controlled experiments using neural style transfer (NST) [14], before evaluating on real data in the next section.
+This allows us to systematically adjust an object’s textural contrast with its background while keeping other factors controlled.
+We begin with all objects sampled from VOC2012 [12] that occupy 5–25% of their image area, to ensure sufficient background context (484 objects total).
+Using a Stable Diffusion-based inpainting model (App.
+C.2), we remove each object and reconstruct its background, creating pairs of separate object-only and background-only images.
+This setup allows precise modifications before reinserting objects into their original scenes.
+For each object, we then create three types of composite images with varying degrees of changes to the object’s shape (without modifying the background):
   1. 1.
@@
   2. 2.
-    Altered: the object undergoes major non-affine geometric transformations (see App. C.2), resulting in shape modification and its texture being squeezed or expanded.
+    Altered: the object undergoes significant non-affine geometric transformations, modifying its shape and distorting its texture (App. C.2).
   3. 3.
-    Mixed: this variant uses pixels from the altered image where the original and altered masks overlap, and pixels from the original object where they do not overlap.
-After one of these three composite image types is created, we use neural style transfer (NST) to apply a texture to the image (similar to experiments in [16]) randomly sampled from Colored Brodatz [1], via an implementation of the NST model of Gatys et al.
-[15] (details in Appendix C.2).
-We illustrate an example of this entire procedure in Fig.
+    Mixed: a hybrid of the original and altered object, using altered pixels where the masks overlap and original pixels elsewhere.
+After creating these image variants, we apply NST (similar to [15]) using textures randomly sampled from Colored Brodatz [1] with the Gatys et al.
+NST model [14] (details in App.
+C.2).
+An example of this entire procedure is in Fig.
 6.
-Figure 6: Example images from each component of the synthetic textural separability dataset creation pipeline.
-We adjust textural separability by performing the style transfer at eight different degrees of severity using different settings for the weighting hyperparameters that control the balance between content preservation and style transfer (details in Appendix C.2), with example styled images shown in Fig.
-7 (more in Appendix.
-A.3).
-In Fig.
-8 left, we validate this scheme by showing that indeed, steadily increasing the NST intensity results in decreasing textural separability.
-Figure 7: Using style transfer with different intensities to adjust the textural separability of an object (highlighted in red), on controlled (upper row) and altered (lower row) versions of the object.
+Figure 6: Example from the NST image creation pipeline.
+To systematically vary textural separability, we apply NST at eight different intensity levels, adjusting the balance between content preservation and style transfer (details in App.
+C.2).
+Fig.
+7 provides examples of this, and Fig.
+8 (left) confirms that increasing NST intensity steadily reduces textural separability as intended.
+Figure 7: Using NST with different intensities to adjust the textural separability of an object (highlighted in red), on controlled (upper row) and altered (lower row) versions of the object.
 Figure 8: Left: Increasing NST intensity results in decreased textural separability. Right: SAM segmentation IoU vs. NST intensity, on the three composite image types.
 In Fig.
-8 right, we show how SAM’s segmentation performance changes with respect to style transfer intensity on the test set, for each of the three types of object transformations.
-We first see that SAM’s performance decreases with lower textural separability/higher NST intensity.
-Second, there is an interesting pattern of clear gaps in segmentation performance between the three object transformation types.
-SAM performs best on the altered objects/foregrounds, which have both distorted shape and texture, compared to the un-altered/controlled objects, with the mixed objects in-between.
-This further demonstrates the effect of textural contrast on segmentation performance, as the altered objects have additional textural contrast to their backgrounds due to the transformations which were applied to their shape, boundary and texture.
+8 right, we analyze how SAM’s performance varies with NST intensity across the three object transformation types.
+First, as textural separability decreases (higher NST intensity), SAM’s performance declines, confirming the impact of textural contrast on segmentation accuracy.
+Second, we observe clear performance gaps between transformation types: SAM performs best on altered objects, which undergo both shape and texture distortion, followed closely by mixed objects, and then performs worst on controlled objects.
+This suggests that additional textural contrast—introduced by shape, boundary, and texture transformations—enhances segmentation performance, further reinforcing the role of textural cues in SFM behavior.
 5.2.2 Experiments on Real Data
 We will now evaluate datasets of real images; iShape (Sec.
-2.3.2) and Plittersdorf [18], both of which were used in the original SAM paper [27], and possess objects with a wide range of textural separability from their surroundings.
+4.3.2) and Plittersdorf [17], both of which were used in the original SAM paper [26], and possess objects with a wide range of textural separability from their surroundings.
 Plittersdorf consists of camera trap video frames of wild deer recorded in a wildlife park.
-These frames often have low contrast objects due to frequent low-light conditions, making it a useful dataset for this analysis.
-Example images and objects from both datasets are shown in Fig.
-9.
-Figure 9: Example images and object masks from Plittersdorf (left group) and iShape (right group).
+These frames sometimes have low-contrast objects due to frequent low-light conditions, making it a useful dataset to analyze.
+Example images and objects from both datasets are shown in App.
+A.4.
 In Fig.
-10, we show how the textural separability (Algorithm 3) of these objects relates to SAM’s segmentation performance on them, with correlation results shown in Table 3.
-Overall, across all datasets we see a fairly strong correlation between textural separability and segmentation performance (average correlation of $\tau=0.49$ and $\rho=0.66$), especially considering the variety of objects and backgrounds, with objects with low separability resulting in especially poor segmentation ($\mathrm{IoU}\sim 0.3$).
-In these cases, SAM was confused by objects close to the object of interest which also have similar textures (see e.g., Fig.
-10, right).
+9, we show how the textural separability (Algorithm 3) of these objects relates to SAM’s performance on them, with correlation results shown for all three SFMs in Table 3.
+Overall, across all datasets, we see a fairly strong correlation between textural separability and segmentation IoU (average corr.
+of $\tau=0.47$ and $\rho=0.62$), especially considering the variety of objects and backgrounds, with objects with low separability typically being noticeably harder to segment.
+In these cases, the SFMs were confused by objects close to the object of interest which also have similar textures.
 iShape
 antenna branch fence hanger log wire Plittersdorf
-SAM Enc. $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
-ViT-H 0.44 0.59 0.50 0.68 0.65 0.85 0.63 0.83 0.33 0.48 0.49 0.67 0.26 0.38
-ViT-B 0.46 0.61 0.56 0.75 0.67 0.86 0.65 0.84 0.36 0.52 0.55 0.73 0.36 0.50
-Table 3: Rank correlation between SAM prediction IoU and object textural separability, on iShape and Plittersdorf.
-Figure 10: SAM prediction IoU vs. object textural separability (Algorithm 3), for iShape (left) and Plittersdorf (right).
-Discussion
-Understanding and Interpreting SAM’s Failure Modes.
-First, we see that SAM’s performance issues on cases of high tree-likeness (Sec.
-2.3) were primarily due to over-segmentation, rather than under-segmentation; the model would frequently generate large false-positive regions in the empty space of densely packed, thin tree-like structures (e.g., Figs.
-3, right and 1, left).
-This could not simply be due to image resolution/downsampling blurring the thin objects, as for structures which are also thin but more sparsely packed, the over-segmentation is reduced (see e.g., Fig.
+Model $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
+SAM, ViT-H [26] 0.44 0.59 0.50 0.68 0.65 0.85 0.63 0.83 0.33 0.48 0.49 0.67 0.26 0.38
+SAM, ViT-B [26] 0.46 0.61 0.56 0.75 0.67 0.86 0.65 0.84 0.36 0.52 0.55 0.73 0.36 0.50
+SAM 2, ViT-L [36] 0.46 0.64 0.56 0.76 0.64 0.83 0.65 0.84 0.25 0.36 0.54 0.71 0.20 0.29
+SAM 2, ViT-B+ [36] 0.44 0.61 0.48 0.67 0.69 0.88 0.64 0.83 0.20 0.28 0.50 0.67 0.12 0.17
+HQ-SAM, ViT-H [24] 0.39 0.54 0.50 0.68 0.69 0.88 0.62 0.80 0.20 0.28 0.47 0.63 0.20 0.28
+HQ-SAM, ViT-B [24] 0.43 0.60 0.50 0.68 0.68 0.86 0.62 0.81 0.23 0.33 0.49 0.66 0.21 0.31
+Average 0.44 0.60 0.52 0.70 0.67 0.86 0.64 0.83 0.26 0.38 0.51 0.68 0.23 0.32
+Table 3: Rank correlation between SFM IoU and object textural separability, on iShape and Plittersdorf.
+Figure 9: Segmentation IoU vs. object textural separability, for iShape (left) shown for SAM, and Plittersdorf (right) for all SFMs.
+6 Robustness to Fine-tuning
+Despite being designed for thin, branching structures, HQ-SAM exhibits performance drops on high tree-likeness objects, similar to SAM and SAM 2 (Table 2, Fig.
+5).
+This suggests that the difficulty SFMs face with such objects persists even after adaptation, a hypothesis we explore here.
+To assess whether fine-tuning could improve SAM’s segmentation of high tree-likeness objects, we selected some $k$ examples from DIS-train using different ranking strategies: Oracle (direct IoU-based selection), Rand (random selection), DoGD / CPR (selecting examples based on high tree-likeness), and Sum (a combination of DoGD and CPR).
+We fine-tuned only SAM’s mask decoder on these, keeping the image encoder frozen to preserve its generalist capabilities (as in [24]); full details are in App.
+C.3.
+Results (Fig.
+10) show that fine-tuning on high tree-likeness examples (DoGD, CPR, or Sum) led to slightly worse validation performance than random selection, likely due to overfitting on these atypical examples, harming generalization.
+Even though fine-tuning improved overall average IoU, the correlation between segmentation performance and tree-likeness remained—both on unseen validation data (middle plots, Fig.
+10) and even on the fine-tuned examples themselves (lower plots).
+These findings indicate that fine-tuning alone is not a viable solution to overcoming SFMs’ performance drop on highly tree-like objects.
+Figure 10: Top: Performance (IoU) of fine-tuned SAM on DIS val vs.
+number of fine-tuning examples $k$ sampled from DIS train, using different selection strategies.
+Lower rows: IoU vs.
+tree-likeness for each $k=1024$ fine-tuned model on DIS val (middle row) and on the fine-tuning set (bottom row).
+7 Discussion
+Understanding SFM Failure Modes.
+We find that SFMs struggle primarily with over-segmentation rather than under-segmentation in high tree-likeness cases (Sec.
+4.3).
+The models often generate large false-positive regions within densely packed, thin structures (e.g., Figs.
+3, right; 1, left).
+This cannot be explained by resolution limitations alone, as sparsely packed thin structures exhibit reduced over-segmentation (Fig.
 3, center rightmost).
-We therefore hypothesize that the root of SAM’s failure on these objects may lie in the highly repetitive yet irregular, dense patterns inherent to tree-like structures in particular: SAM confuses these patterns of shape as the texture of a more regularly-shaped object.
-In the cases of both tree-like and low-textural contrast objects, SAM either has trouble with correctly delineating an object’s shape when faced with a “confusing” textural cue, be it either a dense, irregular shape which appears to instead be a texture, or an object texture that is similar to its surroundings.
-One explanation for this is that SAM’s training set SA-1B [27] possessed few objects with these qualities.
-This seems likely for tree-like objects, given the typically low concavity of objects in SA-1B (Fig.
-6, right in [27]), which were captured in photographic contexts where such objects are uncommon.
-Low-textural contrast objects are similarly uncommon in most photographs, and furthermore, segmenting objects is generally more challenging if they are harder to pick out from their surroundings.
-Applications of Our Findings and Metrics.
-A simple application of our findings and proposed metrics for the tree-likeness and textural separability of objects would be to predict if SAM is expected to perform underwhelmingly on new data, according to if the objects’ measured tree-likeness is high or the textural separability is low.
-Similarly, the general diversity of objects in a segmentation foundation model’s training/fine-tuning dataset could be measured according to the dataset’s distribution of these metrics, which could inform whether additional training data is needed to be acquired/annotated to develop a better generalist model.
+Instead, we hypothesize that SFMs misinterpret dense, irregular patterns as texture, confusing these structures with the appearance of a more uniformly shaped object.
+In the cases of both tree-like and low-textural contrast objects, SFMs struggle to delineate object boundaries when faced with misleading textural cues, be it either a dense, irregular shape that appears to instead be a texture (the former case), or an object texture that is similar to its surroundings (the latter).
+This suggests that SAM’s training set (SA-1B) contained few such objects, as tree-like structures and low-contrast objects are uncommon in typical photographic datasets [26] (for the former, see e.g., the typically low object concavity of SA-1B in Fig.
+6, right in [26]).
+Since HQ-SAM and SAM 2 were derived from SAM/its training set, their similar behavior is unsurprising.
+On HQ-SAM and Fine-tuning.
+Despite being designed and adapted for thin, intricate structures [24], HQ-SAM still exhibits performance drops correlated with tree-likeness (Tables 1, 2), albeit with slight improvements over SAM and SAM 2 (Fig.
+5, left).
+This may stem from HQ-SAM’s adapter-based approach, which leaves SAM’s image encoder and mask decoder frozen, preserving SAM’s limitations.
+Our fine-tuning experiments (Sec.
+6) further suggest that this issue is not easily mitigated, indicating that these structures pose a fundamental segmentation challenge.
 Limitations and Future Work.
-For our findings of the correlations between the object characteristics measured by our metrics and SAM’s performance in segmenting such objects, we attempted to minimize the influence of other confounding factors which could effect segmentation performance by first establishing a baseline with our carefully-controlled synthetic data experiments.
-We then chose real evaluation datasets which covered a wide range of object types according to these metrics in order to gain a better “signal-to-noise ratio” for the studied trends.
-However, it is impossible to mitigate all potential confounding factors which could affect performance in real data, which is why our correlation findings are still noticeable for the real datasets, yet not quite as tight as on the synthetic data.
-We do not evaluate the 2D version of the recently released SAM 2 model due to it likely resulting in similar behavior as SAM 1, along with evidence that oracle predictions of SAM and SAM 2 are typically similar in performance to within $\lesssim 0.1$ IoU [12].
-Another future direction would be to explore the specific component(s) of SAM which can be most attributed to these failures (in a mechanistic sense), which could guide further model development.
-We deem it likely that the susceptible component of SAM is the image encoder, simply because it is the main backbone of the model which extracts features to be used by the lightweight mask decoder, but obtaining a more fine-grained answer would require a careful treatment.
-Conclusion
-In this paper, we quantitatively modeled how the segmentation performance of SAM relates to certain measurable object characteristics: tree-likeness and textural separability.
-We find SAM’s performance to typically be noticeably correlated with these factors, showing the need for further work in understanding and potentially mitigating such “failure modes” of vision foundation models.
+To ensure our findings were robust, we minimized confounding factors by first establishing trends in controlled synthetic experiments before validating them on real datasets spanning diverse object types.
+However, real-world segmentation is inherently complex, making it impossible to eliminate all confounding variables—which may explain why observed correlations were stronger in synthetic data than in real datasets.
+A key future direction is to identify the specific model components responsible for these failures.
+SAM’s image encoder certainly plays a central role, as it serves as the primary feature extractor for the lightweight mask decoder, but a more precise mechanistic analysis could guide architectural improvements to better handle complex structures.
+Conclusion.
+In this paper, we quantitatively modeled how SFM segmentation performance relates to tree-likeness and textural separability.
+Our findings reveal a consistent correlation between these object characteristics and segmentation accuracy, highlighting the need for further research to better understand and mitigate these failure modes—especially since even targeted fine-tuning fails to fully resolve them.
 Acknowledgements
@@
 The retinal blood vessel and road satellite image object masks which we use to generate the synthetic images for the experiments in Sec.
-2.3.1, are from the Retina Blood Vessel [23] and the Road Extraction Challenge data of DeepGlobe18 [8].
+4.3.1, are from the Retina Blood Vessel [22] and the Road Extraction Challenge data of DeepGlobe18 [8].
 For our retinal vessel mask set, we use the vessel masks from the training set of 80 image/mask pairs, and for our road mask set, we randomly sample the same number of masks from the split-merged full dataset, in order to ensure that these two object types appear equally in the synthetic dataset of Sec.
-2.3.1.
+4.3.1.
 A.2 Synthetic Tree-like Object Images
@@
 Object masks shown in the right-most column, with images created using randomly chosen different foreground and background texture combinations in the left columns (before the majority-voting procedure described in Sec.
-2.3.1).
+4.3.1).
+A.4 Example Images from Plittersdorf and iShape
+We show example images from Plittersdorf and iShape in Fig. 13.
+Figure 13: Example images and object masks from Plittersdorf (left group) and iShape (right group).
 Appendix B Hyperparameter/Ablation Studies
@@
 Appendix C Additional Experimental Details
-C.1 SAM Prompting Strategies
-In addition to the SAM prompting details presented in Sec.
-1, all real and style-transferred (NST) images are prompted with (1) a tight bounding-box and (2) several positive and/or negative prompts randomly sampled from either the foreground or background of the object mask, respectively, with respective counts $n_{pos}$ and $n_{neg}$.
+C.1 SFM Prompting Strategies
+In addition to the SFM prompting details presented in Sec.
+3, all real and style-transferred (NST) images are prompted with (1) a tight bounding-box and (2) several positive and/or negative prompts randomly sampled from either the foreground or background of the object mask, respectively, with respective counts $n_{pos}$ and $n_{neg}$.
 For iShape, we use $n_{pos}=n_{neg}=5$; for DIS, we use $n_{pos}=5$ and $n_{neg}=10$ due to the complexity of the objects.
@@
 For the NST experiments (Sec.
-3.2.1), the inpainting model used is Runway’s stable diffusion inpainting pipeline, fp16 variant.
+5.2.1), the inpainting model used is Runway’s stable diffusion inpainting pipeline, fp16 variant.
 The NST model itself is an implementation of [14] based on https://github.com/pytorch/tutorials/blob/main/advanced_source/neural_style_tutorial.py, using typical settings of a standard VGG19 CNN with content layer of conv_4, style layers of conv_i for i = 1, 2, 3, 4, 5, and ImageNet normalization.
@@
 As mentioned in Sec.
-3.2.1, we perform style transfer experiments for a monotonically-increasing range of eight degrees of style transfer intensity.
+5.2.1, we perform style transfer experiments for a monotonically-increasing range of eight degrees of style transfer intensity.
 This is created by defining content and style weights $\lambda_{c}$ and $\lambda_{s}$, respectively for the NST algorithm according to
@@
 shape_transformation,raw_img_arr,raw_msk_arr,raw_msk_arr.copy())
+C.3 Fine-Tuning Experiment Details
+Here we provide all details for the SAM fine-tuning experiments of Section 6.
+The loss function that we use is the sum of the Dice and cross-entropy losses (as implemented by MONAI [5]), optimized via AdamW [29] with a learning rate of $10^{-5}$ and a weight decay strength of $0.1$.
+We use a batch size of $16$ and fine-tune for $40$ epochs on the $k$ (image, mask) examples which were sampled from DIS-train using the given selection strategy (see Section 6).
+All parameters of SAM’s mask decoder are fine-tuned, while the image encoder is frozen.
 Appendix D Additional Results
@@
 Table 11: Linear (Pearson $r$) and nonlinear (Spearman $\rho$, Kendall $\tau$) correlations between CPR and DoGD on all evaluated datasets.
+D.2 Effect of Object Line Thickness on Performance
+In Section 4.1, we hypothesized that the noticeable SAM segmentation performance difference between retinal vessel images and satellite road images, two types of objects with branching structures, is due to the former objects typically having more dense and irregularly-spaced branching structures than the latter.
+However, it is important to consider the possibility that this is simply due to retinal vessels having thinner lines than road objects, which would help evaluate if the proposed tree-likeness metrics (CPR and DoGD) are actually characterizing the true underlying features that drive performance.
+To test this, we performed the same synthetic data experiments as in Section 4.3.1, but applied transformations which thickened object lines, to see how this affected the overall relationship between measured object tree-likeness and segmentation performance.
+To do so, between steps 2 and 3 of the procedure used to generate these images (App.
+A.2), we:
+  1. 1.
+    Converted the object mask $m_{c}$ to its skeleton, effectively resulting in a mask consisting of lines with 1-pixel widths.
+  2. 2.
+    Apply a dilation transformation with a radius of 4 pixels to this skeleton, resulting in each path/line structure in the mask being strictly 9 pixels wide.
+We show the results of this procedure (how it affects the relationship between segmentation IoU and tree-likeness) in Fig.
+15.
+We see that the result of IoU and object tree-likeness being correlated is mostly invariant to the object line thickening transformation.
+However, while the DoGD results are generally maintained because DoGD quantifies the deviation of filling-rate at different scales, we see that CPR experiences some degradation in its correlation with IoU for certain retinal vessel components.
+This is possibly due to the highly dense tree structures, originally thin with varying widths, forced to be similarly thickened and then possibly merged with each other, thereby reducing the overall CPR of the object.
+This may be a limitaiton of CPR in quantifying the tree or net-like geometric structure.
+Fortunately, such special cases where all object pixels fall around a fixed range of object skeletons are rare in realistic scenarios.
+Figure 15: SAM segmentation IoU vs. object tree-likeness (CPR, left; DoGD, right) on the synthetic dataset with dilated object lines.

### v3

@@
 On carefully controlled synthetic experiments and real-world datasets, we show that SFM performance (e.g., SAM, SAM 2, HQ-SAM) noticeably correlates with these factors.
-We link these failures to “textural confusion”, where models misinterpret local structure as global texture, causing over-segmentation or difficulty distinguishing objects from similar backgrounds.
+We attribute these failures to SFMs misinterpreting local structure as global texture, resulting in over-segmentation or difficulty distinguishing objects from similar backgrounds.
 Notably, targeted fine-tuning fails to resolve this issue, indicating a fundamental limitation.
 Our study provides the first quantitative framework for modeling the behavior of SFMs on challenging structures, offering interpretable insights into their segmentation capabilities.
+[^1] Code: https://github.com/mazurowski-lab/SAMFailureMetrics
 1 Introduction
-Segment Anything Model (SAM) [26], the most widely-used segmentation foundation model (SFM), has demonstrated promising zero-shot and fine-tuned segmentation ability across diverse domains, including biomedical imaging [31, 21, 10], remote sensing [6, 37], and others.
-However, certain failure modes for SAM have been found empirically, where SAM produces underwhelming performance on datasets with certain uncommon objects or atypical contexts, such as dense, branching structures (e.g., retinal blood vessels) [38, 11], concealed/low-contrast objects [7, 41, 23, 44], small or irregular objects, and others [23].
-We will show that newer SFMs, including SAM 2 [36] and HQ-SAM [24], are also susceptible to these failure modes, despite some stated improvements over these limitations for SAM—HQ-SAM on thin, branching structures and SAM 2 on low-contrast objects.
-Fine-tuning or adapting these models for such challenging examples using various strategies [30, 33, 16, 24] (or utilizing specialized techniques, e.g., matting algorithms [28]) could mitigate these issues, yet we will show that these strategies do not fully resolve the problems, nor explain the root causes.
+Segment Anything Model (SAM) [sam], the most widely-used segmentation foundation model (SFM), has demonstrated promising zero-shot and fine-tuned segmentation ability across diverse domains, including biomedical imaging [MAZUROWSKI2023102918, huang2024segment, osti_10447851], remote sensing [chen2024rsprompter, ren2024segment], and others.
+However, certain failure modes for SAM have been found empirically, where SAM produces underwhelming performance on datasets with certain uncommon objects or atypical contexts, such as dense, branching structures (e.g., retinal blood vessels) [shi2023generalist, dong_efficient_2024], concealed/low-contrast objects [chen2023sam, tang2023can, ji2024segment, wang2025orderaware], small or irregular objects, and others [ji2024segment].
+We will show that newer SFMs, including SAM 2 [sam2] and HQ-SAM [hqsam], are also susceptible to these failure modes, despite some stated improvements over these limitations for SAM—HQ-SAM on thin, branching structures and SAM 2 on low-contrast objects.
+Fine-tuning or adapting these models for such challenging examples using various strategies [ma2024segment, peng2024sam, gu2024build, hqsam] (or utilizing specialized techniques, e.g., matting algorithms [lin2021real]) could mitigate these issues, yet we will show that such strategies do not fully resolve the problems, nor explain the root causes.
 This underscores the need for a theoretical understanding of why SFMs struggle with such cases, paving the way for better model design and adaptation for new applications.
 Figure 1: SAM’s segmentation performance tends to drop noticeably when the object has high tree-likeness (left, on DIS) or low textural separability (right, on iShape)–even with heavy prompting–which we investigate in this work.
-From left to right: input images, object annotations, and their SAM segmentations.
 Here, we propose that several of these failure modes in SFMs stem from misinterpreting textural cues when distinguishing objects from their surroundings.
 Specifically, SFMs often confuse dense, tree-like structures as large-scale textures, leading to over-segmentation, and struggle with objects that have similar textures to their surroundings.
-To quantify these challenges, we introduce interpretable metrics for object tree-likeness and textural separability.
+To quantify these challenges, we introduce interpretable metrics for object tree-likeness and textural separability, designed to quantify specific aspects of objects that may make them challenging to segment by typical SFM architectures.
 Tree-likeness (Sec.
-4.2) captures structural complexity using spatial contour frequency and the difference in object variability between global and local scales.
+4.2) captures structural complexity with two new metrics, involving spatial contour frequency and the difference in object variability between global and local scales, inspired by how vision transformer encoders and attention patterns in SFMs can be “confused” by tree-like objects.
 Textural separability (Sec.
 5.1) measures how distinguishable an object’s textural features are from those of its background.
-These metrics are designed for broad applicability across datasets and object types and are implemented in a PyTorch-compatible form for efficiency.
+These metrics are designed for broad applicability across datasets and objects and are implemented in a PyTorch-compatible form for efficiency.
 We then investigate how tree-likeness and textural separability affect SFM performance through controlled synthetic experiments (Secs.
@@
 These experiments systematically isolate the influence of object shape and texture while minimizing confounding factors.
-Across all datasets, we find that SFM performance (IoU) noticeably correlates with both tree-likeness and textural separability across diverse object and image types.
+Across all datasets, we find that SFM performance (segmentation IoU) noticeably correlates with both tree-likeness and textural separability across diverse object and image types.
 Notably, HQ-SAM, designed for intricate structures, still exhibits these limitations, highlighting the persistence of these challenges.
@@
   1. 1.
-    We introduce quantitative metrics for object tree-likeness—Contour Pixel Rate (CPR) and Difference of Gini Impurity Deviation (DoGD)—and an object textural separability metric based on neural feature contrast.
+    We introduce quantitative novel metrics for object tree-likeness—Contour Pixel Rate (CPR) and Difference of Gini Impurity Deviation (DoGD)—and an object textural separability metric based on neural feature contrast.
   2. 2.
@@
     We show that these segmentation challenges persist even after fine-tuning, suggesting that SFMs struggle with these object types regardless of adaptation.
-Our findings provide the first quantitative model linking SFM performance to measurable object characteristics.
-Despite the complexity of SFMs and their evaluation datasets, we show that their underperformance on challenging cases can be modeled using our straightforward, interpretable object descriptors.
-We release easy-to-use code for our proposed metrics at https://github.com/mazurowski-lab/SAM-TexturalConfusion-Metrics.
+Our findings provide the first quantitative modeling that links SFM performance to measurable object characteristics.
+Despite the complexity of SFMs and their evaluation datasets, we demonstrate that their underperformance on challenging cases can be effectively modeled using our straightforward and interpretable object descriptors.
 2 Related Work
-In addition to the various “failure modes” and biases of segmentation foundation models discussed in the introduction which motivated this work, various works have also studied the same for general vision models.
-Central to these is the goal of understanding the visual features that networks use to make predictions from images, and whether object shape or texture is more important.
-This is important to study because any unexpected failure of such models will depend on the features used for inference.
+Beyond the original Segment Anything Model (SAM) [sam] and the subsequent SFMs of HQ-SAM [hqsam] and SAM 2 [sam2], other works involving universal, generalist, and/or foundation models or architectures for segmentation include OneFormer [jain2023oneformer], a framework for multi-task universal image segmentation; Segment and Caption Anything (SCA), a method equipping SAM with caption generation [huang2024sca]; Semantic-SAM [li2024semanticsam], for segmenting with controllable granularity between objects; Open-Vocabulary SAM [yuan2024open], a framework unifying SAM and CLIP [clip]; and SEEM [zou2023seem], a less commonly-used multimodal foundation model developed concurrently with SAM.
+Previous work on biases and failure modes in computer vision models for tasks such as segmentation has primarily focused on understanding whether networks rely more on object shape or texture for making visual predictions.
 Zhang et al.
-[49] posited that objects within images can be described by three characteristics: shape, texture, and the composition of textures within shapes, and analyzed how these characteristics may affect the challenge of segmenting such objects.
-It is well known that convolutional neural networks’ (CNNs’) predictions are typically biased towards the texture of objects rather than their shape, for both classification [15, 40, 20, 2] and segmentation [49], resulting in curious examples where network predictions are based on texture even when the shape clearly defines a different object.
-For example, [27] found that when a CNN is presented with images of objects with only their shape visible, not texture (silhouettes), performance was greatly worsened.
-It has also been found that the shape-learning behavior of neural networks can differ noticeably depending on the particular training dataset and objective [20, 19], motivating us to study the specific behavior of SAM (given its unique training set) and other similar SFMs in detail.
-Interestingly, there has been some evidence that vision transformers (like SAM’s image encoder) may instead be biased towards shape, rather than texture for classification tasks [42].
-Here, we study segmentation rather than classification, where it is unclear if those results extend, as our results generally point to SAM focusing on texture over shape (to the point of even confusing complex, locally varying shape itself as texture), which aligns with the preliminary findings of [48].
+[zhang2024convolutional] proposed a framework describing objects by shape, texture, and texture composition within shapes, highlighting how these characteristics impact segmentation tasks.
+Extensive literature shows that convolutional neural networks (CNNs) tend to be biased toward texture over shape, affecting performance in both classification [geirhosimagenet2019, subramanian2024spatial, hermann2020origins, baker2018deep] and segmentation [zhang2024convolutional] tasks.
+Notably, CNN performance significantly drops when tested on shape-only images (silhouettes), underscoring texture bias [kubilius2016deep].
+Moreover, neural networks’ shape-learning behaviors vary based on the training dataset and objectives [hermann2020origins, hermann2020shapes].
+This variation motivates our detailed investigation into the specific behaviors of SFMs such as the Segment Anything Model (SAM), given their unique training characteristics.
+While vision transformers (such as SAM’s image encoder architecture) have demonstrated a preference for shape in classification tasks [tuli2021convolutional], it remains uncertain if this extends to segmentation.
+Our results suggest SAM predominantly leverages texture cues, sometimes even interpreting complex shapes as textures, corroborating prior studies [zhang2023understanding].
 3 Methods: SFM Usage and Prompting
-We experiment with the ViT-H (default, if unstated) and ViT-B standard pre-trained SAM models [26], the ViT-L and ViT-B+ variants of SAM 2 [36], and the ViT-H and ViT-B variants of HQ-SAM [24], a SAM adaptation designed particularly for thin structures.
+We experiment with the ViT-H (default, if unstated) and ViT-B standard pre-trained SAM models [sam] (labeled “SAM-H” and “SAM-B”), the ViT-L and ViT-B+ variants of SAM 2 [sam2] (“SAM 2-L” and “SAM 2-B+”), and the ViT-H and ViT-B variants of HQ-SAM [hqsam] (“HQ-SAM-H” and “HQ-SAM-B“), a SAM adaptation designed particularly for thin structures.
 As is the default for SAM, all input images are resized to a resolution of $1024\times 1024$, and normalized to $[0,255]$.
 All images that possess instance segmentations for multiple objects will be evaluated by the given SFM segmenting each of the images’ objects one at a time (guided by clear prompting described in the next paragraph), with all other mask pixels set as background.
-In all experiments, we use the oracle-chosen mask out of the given SFM’s three outputted predictions.
+In all experiments, we use the oracle-chosen prediction from the given SFM.
 For all experiments on the synthetic images of Sec.
@@
 4.1 Motivation
-We first wished to understand the challenges for SFMs in segmenting certain objects with dense, tree-like structures, and then quantify the features driving these challenges, following our observation of a trend of SAM having difficulty with such objects, including retinal blood vessels [35, 30] or satellite road images [13, 45] (see e.g., Fig.
+We first wished to understand the challenges for SFMs in segmenting certain objects with dense, tree-like structures, and then quantify the features driving these challenges, following our observation of a trend of SAM having difficulty with such objects, including retinal blood vessels [qiu2023learnable, ma2024segment] or satellite road images [feng2024road, xu2023leveraging] (see e.g., Fig.
 2).
 Interestingly, while both of these object types have branching features, a characteristic which naively could relate to the SFM’s “failure”, SAM’s performance on retinal vessel images is noticeably worse (avg.
-$\mathrm{IoU}\simeq 0.05$ [35]) compared to on satellite road structures (avg.
-$\mathrm{IoU}\simeq 0.2$ [13]), which was reproduced in our own experiments (Fig.
+$\mathrm{IoU}\simeq 0.05$ [qiu2023learnable]) compared to on satellite road structures (avg.
+$\mathrm{IoU}\simeq 0.2$ [feng2024road]), which was reproduced in our own experiments (Fig.
 4) [^2] This could potentially simply be due to retinal vessels being thinner than road structures, but our experiments do not support this (App.
-D.2)..
+D.4)..
 Figure 2: Example retinal blood vessel (top) and satellite road (bottom) images and accompanying object segmentation masks.
-We hypothesize that such objects become more challenging to segment when their branching structures are dense and irregularly-spaced, which we refer to as “tree-like”.
-We will show that SFMs consider such structures as textures rather than shapes, resulting in significant over-segmentation.
-In the following section, we propose how to characterize the tree-likeness of objects using two new quantitative metrics: CPR (Contour Pixel Rate) and DoGD (Difference of Gini Impurity Deviation).
+We hypothesize that objects with dense and irregularly-spaced branching structures–which we refer to as ”tree-like”–pose fundamental challenges to SFMs because their structural properties create a mismatch with patch-based attention mechanisms.
+When a ViT [dosovitskiy2020image] divides an object into fixed-size patches, tree-like structures present a critical issue: their boundary pixels are distributed throughout the object rather than concentrated at the perimeter.
+Consequently, most patches contain predominantly boundary information, leaving insufficient object-internal features to build coherent representations.
+In the following section, we will formalize this via our proposed Contour Pixel Rate (CPR) metric.
+SFM self-attention maps for tree-like objects for SAM then have more fragmented, locally-focused attention patterns that respond to local edges independently rather than forming coherent global object representations, which we show empirically (via the spatial autocorrelation of attention maps) in App.
+D.2.2 with examples shown in App.
+D.1.
+Moreover, we empirically show that typical object CPR increases with a model’s effective patch size (App.
+D.2.1).
+Beyond boundary density, tree-like structures also exhibit dramatic appearance variation across different scales: thick trunks, medium branches, and fine tendrils create inconsistent local patterns.
+We capture this through our Distance of Gini impurity Deviation (DoGD) metric, proposed in the following section, which measures how object pixel distribution varies across scales.
+High DoGD values indicate scale-dependent spatial inconsistency, preventing stable object representations.
+The combination of high CPR and high DoGD explains why SFMs fail on tree-like structures: boundary-dominated patches (high CPR) produce fragmented attention lacking semantic coherence, while scale-dependent variations (high DoGD) prevent consistent object-level representations.
+This causes SFMs to interpret tree-like objects as textures–collections of local edges without global structure–leading to over-segmentation where unified objects are split into multiple regions.
+Our metrics thus provide a quantitative framework for predicting SFM failures on complex branching structures.
 4.2 Quantifying Tree-Like Structures
@@
 We first propose to measure the degree of tree-like structure of the object $m$ according to the percentage of the object’s pixels which lie on it’s contour, which are defined as follows.
-Definition 1 (Contour Pixels).
-Given some mask $m$, a pixel $m_{ij}$ is a contour pixel if $m_{ij}=1$ and there exists a different pixel $m_{kl}$ such that $m_{kl}\neq m_{ij}$ and $||(k,l)-(i,j)||_{1}<R$, given some small contour width threshold $R>0$.
+Definition 4.1 (Contour Pixels).
+Given some mask $m$, a pixel $m_{ij}$ is a contour pixel if $m_{ij}=1$ and there exists a different pixel $m_{kl}$ such that $m_{kl}\neq m_{ij}$ and $(k,l)-(i,j)||_{1}<R$, given some smallthreshold $R>0$.
 In other words, contour pixels have at least one pixel of a different class within a small neighborhood.
@@
 $\mathrm{CPR}(m):=\frac{|C|}{|F|}.$ (1)
-Intuitively, tree-like objects will have a higher percentage of contour pixels, resulting in higher CPRs.
+The more tree-like an object is, the higher the percentage of its pixels that are contour pixels, resulting in a higher CPR.
 We demonstrate fast computation of CPR in full detail in Algorithm 1, via vectorized PyTorch-like pseudocode.
@@
 $\mathrm{Gini}(m;k,h_{0},w_{0}):=$ (3)
-$1-[p(\mathcal{W}^{k}_{h_{0},w_{0}}(m))]^{2}-[1-p(\mathcal{W}^{k}_{h_{0},w_{0}}(m))]^{2},$
-[…]
+$1-[p(\mathcal{W}^{k}_{h_{0},w_{0}}(m))]^{2}-[1-p(\mathcal{W}^{k}_{h_{0},w_{0}}(m))]^{2}$
+where we write $p:=p_{1}$.
+The Gini impurity measures the degree of uncertainty (ranging from 0 to $0.5$) of whether an object is present in the given window.
+For example, mask windows containing pixels of mostly one class will have $\mathrm{Gini}\simeq 0$, while having similar pixel amounts of both classes will result in $\mathrm{Gini}\simeq 0.5$ [^3] Following common decision tree practices, we use Gini impurity over entropy because it is a symmetric and faster alternative [breiman1984classification]..
 Next, we compute the variability of object presence at a given scale/window size across the entire mask by sampling all possible $k\times k$ windows with anchors $(h_{0},w_{0})$, and computing the standard deviation of the Gini impurity across these windows, as
@@
 return gini_std[a] - gini_std[b]
-Intuitively, objects with significant tree-like or fractal-like structure will exhibit relatively large values of $\sigma^{\mathrm{Gini}}_{b}(m)$ due to high variability in pixel composition at small scales (frequently alternating between areas of mixed classes and areas of a single class), yet small $\sigma^{\mathrm{Gini}}_{a}(m)$ due to structures with high uniformity and/or repetitions at large scales, altogether increasing the DoGD.
-We perform all experiments with the hyperparameters for CPR and DoGD set to $R=5$, $a=127$, and $b=3$, which we found via grid search for the values that resulted in the Kendall’s $\tau$ between IoU and DoGD with the lowest $p$-value on the held-out DIS training set using ViT-H SAM.
-We show results using a wide range of other values for these hyperparameters in Appendix B.1, where our findings were consistent for most other settings.
-Moreover, we note that CPR and DoGD are correlated with each other (Pearson $|r|=0.83$ on average; Appendix D.1), showing their consistency in quantifying different aspects of tree-likeness.
+Objects with significant tree-like or fractal-like structure will exhibit relatively large values of $\sigma^{\mathrm{Gini}}_{b}(m)$ due to high variability in pixel composition at small scales (frequently alternating between areas of mixed classes and areas of a single class), yet small $\sigma^{\mathrm{Gini}}_{a}(m)$ due to structures with high uniformity and/or repetitions at large scales, altogether increasing the DoGD.
+We perform all experiments with the hyperparameters for CPR and DoGD set to $R=5$, $a=127$, and $b=3$, which we found via grid search for the values that resulted in the Kendall’s $\tau$ between IoU and DoGD with the lowest $p$-value on the held-out DIS training set using SAM-H.
+We show results using a wide range of other values for these hyperparameters in App.
+B.1, where our findings were consistent for most other settings.
+Moreover, we note that CPR and DoGD are noticeably correlated (Pearson $|r|=0.83$ on average; App.
+D.3), showing their consistency in quantifying different aspects of tree-likeness.
 4.3 The Relationship between Object Tree-likeness and Segmentation Performance
@@
 In this section, we will first probe the effect of object tree-likeness on the performance of SFMs by testing them on synthetic images that solely possess objects of varying tree-likeness, with different independently chosen, uniform foreground and background textures.
-These objects are contiguous components sampled from retinal blood vessel and satellite road masks; the full procedure of generating these images is detailed in Appendix A.2.
+These objects are contiguous components sampled from retinal blood vessel and satellite road masks; the full procedure of generating these images is detailed in App.
+A.2.
 Example generated images, masks, and SAM predictions for them are shown in Fig.
@@
 As shown (also in Fig.
-4), the objects cover a wide range of tree-likeness as measured by these quantities.
+4, left), the objects cover a wide range of tree-likeness as measured by these quantities.
 In order to mitigate any confounding effects on a given SFM’s performance due to an object’s textural contrast from its surroundings (which we study in Sec.
@@
 In Fig.
-4 we show the relationship between an object’s tree-likeness (via CPR or DoGD) with SAM’s performance (IoU) in segmenting the object, for all generated synthetic objects.
-We quantify the strength of this relation via non-linear correlation, measured by Kendall’s tau ($\tau$) [25] and Spearman’s rho ($\rho$) [39], for all three SFMs, shown in Table 1.
+4 left, we show the relationship between an object’s tree-likeness (via CPR or DoGD) with SAM’s performance (IoU) in segmenting the object, for all generated synthetic objects.
+We quantify the strength of this relation via non-linear correlation, measured by Kendall’s tau ($\tau$) [kendall1938new] and Spearman’s rho ($\rho$) [spearman_proof_1904], for all three SFMs, shown in the table to the right of the plots.
 We see that object tree-likeness is quite predictive of an SFM’s performance, with average absolute correlations of $|\tau|=0.75$ and $|\rho|=0.92$ for CPR and $|\tau|=0.67$ and $|\rho|=0.84$ for DoGD; i.e., more prevalent dense tree-like structures corresponding to worse performance.
+In the following section, we will show that this finding is also present for real data, which is subject to noise from a variety of uncontrollable factors.
+Refer to captionRefer to caption
 CPR DoGD
 Model $\tau$ $\rho$ $\tau$ $\rho$
-SAM, ViT-H [26] -0.77 -0.93 0.61 0.80
-SAM, ViT-B [26] -0.75 -0.93 0.66 0.84
-SAM 2, ViT-L [36] -0.73 -0.91 0.67 0.85
-SAM 2, ViT-B+ [36] -0.71 -0.90 0.69 0.86
-HQ-SAM, ViT-H [24] -0.74 -0.92 0.70 0.87
-HQ-SAM, ViT-B [24] -0.81 -0.95 0.66 0.84
+SAM-H -0.77 -0.93 0.61 0.80
+SAM-B -0.75 -0.93 0.66 0.84
+SAM 2-L -0.73 -0.91 0.67 0.85
+SAM 2-B+ -0.71 -0.90 0.69 0.86
+HQ-SAM-H -0.74 -0.92 0.70 0.87
+HQ-SAM-B -0.81 -0.95 0.66 0.84
 Average -0.75 -0.92 0.67 0.84
-Table 1: Rank correlation between SFM IoU and object tree-likeness (CPR and DoGD), on the synthetic dataset.
-Figure 4: SAM IoU vs. object tree-likeness (CPR, left; DoGD, right) on the synthetic dataset.
-While this relationship is certainly strong, it is still on synthetic, controlled data.
-In the following section, we will show that this finding is also present for real data which is subject to noise from a variety of uncontrollable factors.
+Figure 4: Left: SFM prediction IoU vs. object tree-likeness (CPR and DoGD), on the synthetic dataset, shown for SAM-H. Right: Rank correlations between IoU and tree-likeness for all SFMs.
 4.3.2 Experiments on Real Data
-We will now perform the same analysis on two real datasets which contain objects which cover a spectrum of tree-likeness, DIS5k and iShape.
-DIS5k [34] (or “DIS”) is a dataset containing extremely detailed annotations of objects with varying degrees of hollowness, and both regular and irregular tree-like structures.
-All DIS experiments will be reported on its validation set unless otherwise stated.
+We will now perform the same analysis on three real datasets which contain objects which cover a spectrum of tree-likeness, DIS5k, MOSE, and iShape.
+DIS5k [DIS] (or “DIS”) is a dataset containing extremely detailed annotations of objects with varying degrees of hollowness, and both regular and irregular tree-like structures.
+All DIS experiments will be reported on its validation set.
 We show example images with objects with varying degrees of tree-likeness (by CPR) in Fig.
 1 left.
-iShape [46] consists of six sub-datasets of real and realistic-appearing synthetic images for instance segmentation of different objects: antenna, branches, fences, logs, hangers, and wires (see e.g.
+We additionally performed experiments on the MOSE dataset [MOSE] of videos of objects in complex, cluttered, and/or camouflaged scenes, randomly sampled 4 frames with non-empty labels from each of the 1,507 videos in the training set, resulting in $\sim$6,000 frames with $\sim$13,000 total objects.
+iShape [yang2021ishape] consists of six sub-datasets of real and realistic-appearing synthetic images for instance segmentation of different objects: antenna, branches, fences, logs, hangers, and wires (see e.g.
 Fig.
 13).
-We analyze these classes individually to mitigate potential confounding/noise factors due to inter-class variations; we do not do this for DIS due to the larger number of classes which are much more fine-grained in their differences, additionally because there are few images per class for DIS.
+We analyze these classes individually to mitigate potential confounding/noise factors due to inter-class variations.
 In Fig.
-5 we show how SAM’s performance (IoU) on these images relates to the tree-likeness of the objects which it is segmenting, alongside the other SFMs (SAM 2 and HQ-SAM) for comparison for DIS, with accompanying quantitative correlation results for all three SFMs shown in Table 2.
-In order to reduce the noise incurred by the large variety of segmented objects in the dataset and nuisance confounding factors in the images, we analyze results with aggregated objects: we cluster groups of $5$ objects/images with similar IoU and tree-likeness (CPR or DoGD) into single datapoints of the average value of these metrics, and similar for the IoU vs.
+5 left, we show how SAM’s performance (IoU) on these images relates to the tree-likeness of the objects which it is segmenting, alongside the other SFMs (SAM 2 and HQ-SAM) for comparison for DIS, with accompanying quantitative correlation results for all three SFMs shown in the table to the right of the plots.
+The plots for MOSE results are provided in App.
+D.7 due to space constraints.
+In order to reduce the noise incurred by the large variety of segmented objects in the dataset and nuisance confounding factors in the images, we analyze results with aggregated objects: we cluster groups of $5$ objects/images ($20$ for MOSE due to its size) with similar IoU and tree-likeness (CPR or DoGD) into single datapoints of the average value of these metrics, and similar for the IoU vs.
 textural separability experiments of Sec.
 5.
+Refer to captionRefer to captionRefer to captionRefer to caption
 iShape
-DIS antenna branch fence hanger log wire
-Model $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
-CPR: SAM, ViT-H [26] -0.59 -0.76 -0.13 -0.21 -0.61 -0.81 -0.61 -0.81 -0.25 -0.37 -0.60 -0.77 -0.06 -0.10
-SAM, ViT-B [26] -0.63 -0.81 -0.18 -0.27 -0.63 -0.82 -0.63 -0.83 -0.44 -0.62 -0.62 -0.80 -0.12 -0.17
-SAM 2, ViT-L [36] -0.41 -0.57 -0.30 -0.43 -0.64 -0.84 -0.61 -0.81 -0.48 -0.66 -0.67 -0.85 -0.14 -0.21
-SAM 2, ViT-B+ [36] -0.38 -0.52 -0.16 -0.23 -0.52 -0.71 -0.57 -0.78 -0.42 -0.59 -0.60 -0.77 -0.12 -0.18
-HQ-SAM, ViT-H [24] -0.52 -0.69 -0.23 -0.33 -0.77 -0.93 -0.79 -0.95 -0.51 -0.70 -0.69 -0.85 -0.22 -0.33
-HQ-SAM, ViT-B [24] -0.62 -0.81 -0.25 -0.34 -0.75 -0.92 -0.83 -0.96 -0.50 -0.69 -0.71 -0.87 -0.24 -0.35
-Average -0.53 -0.69 -0.21 -0.30 -0.65 -0.84 -0.67 -0.86 -0.43 -0.61 -0.65 -0.82 -0.15 -0.22
-DoGD: SAM, ViT-H [26] 0.58 0.77 0.23 0.30 0.52 0.70 0.46 0.65 0.52 0.72 0.30 0.46 0.08 0.11
-SAM, ViT-B [26] 0.45 0.64 0.12 0.14 0.51 0.69 0.45 0.64 0.50 0.70 0.13 0.22 0.06 0.08
-SAM 2, ViT-L [36] 0.44 0.60 0.42 0.59 0.48 0.65 0.49 0.68 -0.05 -0.08 0.56 0.75 -0.03 -0.05
-SAM 2, ViT-B+ [36] 0.26 0.39 0.40 0.56 0.29 0.40 0.30 0.44 -0.06 -0.10 0.42 0.58 -0.08 -0.12
-HQ-SAM, ViT-H [24] 0.48 0.66 0.53 0.72 0.58 0.72 0.48 0.67 -0.31 -0.45 0.57 0.75 -0.13 -0.19
-HQ-SAM, ViT-B [24] 0.50 0.69 0.67 0.85 0.59 0.73 0.57 0.77 -0.25 -0.34 0.53 0.71 -0.04 -0.06
-Average 0.45 0.63 0.40 0.53 0.50 0.65 0.46 0.64 0.06 0.08 0.42 0.58 -0.02 -0.04
-Table 2: Rank correlation between SFM IoU and object tree-likeness (CPR, upper table and DoGD, lower table), on DIS and iShape.
-Figure 5: Segmentation IoU vs. object tree-likeness (CPR, top; DoGD, bottom) on DIS (left) for all three SFMs and iShape (right) shown for SAM.
+DIS MOSE antenna branch fence hanger log wire
+Model $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
+CPR:
+SAM-H -0.59 -0.76 -0.80 -0.95 -0.13 -0.21 -0.61 -0.81 -0.61 -0.81 -0.25 -0.37 -0.60 -0.77 -0.06 -0.10
+SAM-B -0.63 -0.81 -0.76 -0.93 -0.18 -0.27 -0.63 -0.82 -0.63 -0.83 -0.44 -0.62 -0.62 -0.80 -0.12 -0.17
+SAM 2-L -0.41 -0.57 -0.84 -0.96 -0.30 -0.43 -0.64 -0.84 -0.61 -0.81 -0.48 -0.66 -0.67 -0.85 -0.14 -0.21
+SAM 2-B+ -0.38 -0.52 -0.81 -0.95 -0.16 -0.23 -0.52 -0.71 -0.57 -0.78 -0.42 -0.59 -0.60 -0.77 -0.12 -0.18
+HQ-SAM-H -0.52 -0.69 -0.79 -0.94 -0.23 -0.33 -0.77 -0.93 -0.79 -0.95 -0.51 -0.70 -0.69 -0.85 -0.22 -0.33
+HQ-SAM-B -0.62 -0.81 -0.76 -0.92 -0.25 -0.34 -0.75 -0.92 -0.83 -0.96 -0.50 -0.69 -0.71 -0.87 -0.24 -0.35
+Average -0.53 -0.69 -0.79 -0.94 -0.21 -0.30 -0.65 -0.84 -0.67 -0.86 -0.43 -0.61 -0.65 -0.82 -0.15 -0.22
+DoGD:
+SAM-H 0.58 0.77 0.43 0.60 0.23 0.30 0.52 0.70 0.46 0.65 0.52 0.72 0.30 0.46 0.08 0.11
+SAM-B 0.45 0.64 0.27 0.39 0.12 0.14 0.51 0.69 0.45 0.64 0.50 0.70 0.13 0.22 0.06 0.08
+SAM 2-L 0.44 0.60 0.52 0.70 0.42 0.59 0.48 0.65 0.49 0.68 -0.05 -0.08 0.56 0.75 -0.03 -0.05
+SAM 2-B+ 0.26 0.39 0.47 0.66 0.40 0.56 0.29 0.40 0.30 0.44 -0.06 -0.10 0.42 0.58 -0.08 -0.12
+HQ-SAM-H 0.48 0.66 0.41 0.58 0.53 0.72 0.58 0.72 0.48 0.67 -0.31 -0.45 0.57 0.75 -0.13 -0.19
+HQ-SAM-B 0.50 0.69 0.33 0.48 0.67 0.85 0.59 0.73 0.57 0.77 -0.25 -0.34 0.53 0.71 -0.04 -0.06
+Average 0.45 0.63 0.41 0.57 0.40 0.53 0.50 0.65 0.46 0.64 0.06 0.08 0.42 0.58 -0.02 -0.04
+Figure 5: Left: Segmentation IoU vs.
+object tree-likeness (CPR, top; DoGD, bottom) on DIS (left) for all three SFMs and iShape (right) shown for SAM-H.
+Right: rank correlations between IoU and tree-likeness for all SFMs.
 We see that despite the many potential confounding factors in real data, there is still a clear correlation between object tree-likeness as measured by the proposed metrics and SFM segmentation performance.
-In particular, we see average absolute correlations of $|\tau|=0.59$ and $|\rho|=0.76$ for CPR and $|\tau|=0.38$ and $|\rho|=0.51$ for DoGD—excluding the antenna and wire objects of iShape which had outlying correlations likely because those two object classes cover only small ranges of tree-likeness as opposed to the other iShape classes (Fig.
-5), such that noise from other confounding factors obscures any dependence of performance on tree-likeness.
+In particular, we see average absolute correlations of $|\tau|=0.62$ and $|\rho|=0.79$ for CPR (particularly high for MOSE) and $|\tau|=0.38$ and $|\rho|=0.53$ for DoGD—excluding the antenna and wire objects of iShape which had outlying correlations likely because those two object classes cover only small ranges of tree-likeness as opposed to the other iShape classes (Fig.
+5, left), such that noise from other confounding factors obscures any dependence of performance on tree-likeness.
 5 The Challenge of Textural Separability
@@
 We will define the textural contrast or separability between some object mask and its surroundings by how easily their textures can be distinguished from one another.
-Motivated by findings that early layers of classification-pretrained CNNs primarily capture low-level features involving edges and textures [47], we characterize an image’s textures using the first convolutional layer of a ResNet-18 [18] pretrained on ImageNet [9].
+Motivated by findings that early layers of classification-pretrained CNNs primarily capture low-level features involving edges and textures [zeiler2014visualizing], we characterize an image’s textures using the first convolutional layer of a ResNet-18 [he2016resnet] pretrained on ImageNet [deng2009imagenet].
 Denoted by $f_{1}:\mathbb{R}^{C\times H\times W}\rightarrow\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$, this outputs a textural feature map corresponding to all $7\times 7$ windows in the image.
 We then measure the textural separability of an object according to if a simple classifier $g$ can be trained to discriminate between (a) the activations for the pixels of the object foreground and (b) the activations right outside of the object boundary.
-We define this explicitly in Algorithm 3, where $\mathrm{Dilate}$ and $\mathrm{Disk}$ refer to scikit-image.morphology functions [43].
-We use simple logistic regression for $g$ with inverse regularization parameter $C=2$ [32].
-We also evaluate using other hyperparameter settings, as well as a random forest classifier for $g$ instead, in Appendix B.2, where we found similar results.
+We define this in Algorithm 3, where $\mathrm{Dilate}$ and $\mathrm{Disk}$ refer to scikit-image.morphology functions [scikit-image].
+We use simple logistic regression for $g$ with regularization parameter $C=2$ [scikit-learn].
+We also evaluate other hyperparameter settings, as well as a random forest classifier for $g$, in App.
+B.2, where we found similar results.
 Algorithm 3 Textural separability of an object.
@@
 6:  Train $g$ to classify between activations in $h_{\mathrm{obj}}$ vs. $h_{\mathrm{bdry}}$.
-7:  return  Training classification accuracy of $g$.
+7:  Return Test classification accuracy of $g$.
 5.2 The Effect of Textural Separability on Segmentation Performance
 5.2.1 Style Transfer for Controlled Textural Contrast
-To isolate the effects of textural separability from object shape, we first conduct controlled experiments using neural style transfer (NST) [14], before evaluating on real data in the next section.
+To isolate the effects of textural separability from object shape, we first conduct controlled experiments using neural style transfer (NST) [gatys2016image], before evaluating on real data in the next section.
 This allows us to systematically adjust an object’s textural contrast with its background while keeping other factors controlled.
-We begin with all objects sampled from VOC2012 [12] that occupy 5–25% of their image area, to ensure sufficient background context (484 objects total).
+We begin with all objects sampled from VOC2012 [pascal-voc-2012] that occupy 5–25% of their image area, to ensure sufficient background context (484 objects total).
 Using a Stable Diffusion-based inpainting model (App.
@@
 This setup allows precise modifications before reinserting objects into their original scenes.
-For each object, we then create three types of composite images with varying degrees of changes to the object’s shape (without modifying the background):
+We then create three types of composite images with varying degrees of changes to object shape for each object (without modifying the background):
   1. 1.
@@
     Mixed: a hybrid of the original and altered object, using altered pixels where the masks overlap and original pixels elsewhere.
-After creating these image variants, we apply NST (similar to [15]) using textures randomly sampled from Colored Brodatz [1] with the Gatys et al.
-NST model [14] (details in App.
+After creating these image variants, we apply NST (similar to [geirhosimagenet2019]) using textures randomly sampled from Colored Brodatz [abdelmounaime2013newMBT] with the Gatys et al.
+NST model [gatys2016image] (details in App.
 C.2).
@@
 In Fig.
-8 right, we analyze how SAM’s performance varies with NST intensity across the three object transformation types.
+8 right, we analyze how SAM’s performance varies with NST intensity across the three object transformation types (we found similar results for the other SFMs, shown in App.
+D.5).
 First, as textural separability decreases (higher NST intensity), SAM’s performance declines, confirming the impact of textural contrast on segmentation accuracy.
 Second, we observe clear performance gaps between transformation types: SAM performs best on altered objects, which undergo both shape and texture distortion, followed closely by mixed objects, and then performs worst on controlled objects.
-This suggests that additional textural contrast—introduced by shape, boundary, and texture transformations—enhances segmentation performance, further reinforcing the role of textural cues in SFM behavior.
+This suggests that additional textural contrast–introduced by shape, boundary, and texture transformations–enhances segmentation performance, further reinforcing the role of textural cues in SFM behavior.
 5.2.2 Experiments on Real Data
-We will now evaluate datasets of real images; iShape (Sec.
-4.3.2) and Plittersdorf [17], both of which were used in the original SAM paper [26], and possess objects with a wide range of textural separability from their surroundings.
+We will now evaluate datasets of real images: iShape, MOSE (Sec.
+4.3.2), and Plittersdorf [haucke2022socrates]—of which iShape and Plittersdorf were used in the original SAM paper [sam]—which all possess objects with a wide range of textural separability from their surroundings.
 Plittersdorf consists of camera trap video frames of wild deer recorded in a wildlife park.
@@
 In Fig.
-9, we show how the textural separability (Algorithm 3) of these objects relates to SAM’s performance on them, with correlation results shown for all three SFMs in Table 3.
+9 left, we show how the textural separability (Algorithm 3) of these objects relates to SAM’s performance on them, with correlation results shown for all three SFMs to the right of the plots (MOSE plots shown in App.
+D.7).
 Overall, across all datasets, we see a fairly strong correlation between textural separability and segmentation IoU (average corr.
-of $\tau=0.47$ and $\rho=0.62$), especially considering the variety of objects and backgrounds, with objects with low separability typically being noticeably harder to segment.
+of $\tau=0.49$ and $\rho=0.65$), especially considering the variety of objects and backgrounds, with objects with low separability typically being noticeably harder to segment.
+This relationship was especially strong for the MOSE dataset, with average $\tau=0.65$ and $\rho=0.82$.
 In these cases, the SFMs were confused by objects close to the object of interest which also have similar textures.
+Refer to captionRefer to caption
 iShape
-antenna branch fence hanger log wire Plittersdorf
-Model $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
-SAM, ViT-H [26] 0.44 0.59 0.50 0.68 0.65 0.85 0.63 0.83 0.33 0.48 0.49 0.67 0.26 0.38
-SAM, ViT-B [26] 0.46 0.61 0.56 0.75 0.67 0.86 0.65 0.84 0.36 0.52 0.55 0.73 0.36 0.50
-SAM 2, ViT-L [36] 0.46 0.64 0.56 0.76 0.64 0.83 0.65 0.84 0.25 0.36 0.54 0.71 0.20 0.29
-SAM 2, ViT-B+ [36] 0.44 0.61 0.48 0.67 0.69 0.88 0.64 0.83 0.20 0.28 0.50 0.67 0.12 0.17
-HQ-SAM, ViT-H [24] 0.39 0.54 0.50 0.68 0.69 0.88 0.62 0.80 0.20 0.28 0.47 0.63 0.20 0.28
-HQ-SAM, ViT-B [24] 0.43 0.60 0.50 0.68 0.68 0.86 0.62 0.81 0.23 0.33 0.49 0.66 0.21 0.31
-Average 0.44 0.60 0.52 0.70 0.67 0.86 0.64 0.83 0.26 0.38 0.51 0.68 0.23 0.32
-Table 3: Rank correlation between SFM IoU and object textural separability, on iShape and Plittersdorf.
-Figure 9: Segmentation IoU vs. object textural separability, for iShape (left) shown for SAM, and Plittersdorf (right) for all SFMs.
+antenna branch fence hanger log wire Plitters. MOSE
+Model $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
+SAM-H 0.44 0.59 0.50 0.68 0.65 0.85 0.63 0.83 0.33 0.48 0.49 0.67 0.26 0.38 0.65 0.83
+SAM-B 0.46 0.61 0.56 0.75 0.67 0.86 0.65 0.84 0.36 0.52 0.55 0.73 0.36 0.50 0.74 0.90
+SAM 2-L 0.46 0.64 0.56 0.76 0.64 0.83 0.65 0.84 0.25 0.36 0.54 0.71 0.20 0.29 0.58 0.76
+SAM 2-B+ 0.44 0.61 0.48 0.67 0.69 0.88 0.64 0.83 0.20 0.28 0.50 0.67 0.12 0.17 0.58 0.75
+HQ-SAM-H 0.39 0.54 0.50 0.68 0.69 0.88 0.62 0.80 0.20 0.28 0.47 0.63 0.20 0.28 0.64 0.82
+HQ-SAM-B 0.43 0.60 0.50 0.68 0.68 0.86 0.62 0.81 0.23 0.33 0.49 0.66 0.21 0.31 0.72 0.88
+Average 0.44 0.60 0.52 0.70 0.67 0.86 0.64 0.83 0.26 0.38 0.51 0.68 0.23 0.32 0.65 0.82
+Figure 9: Left: Segmentation IoU vs.
+object textural separability, for iShape (left) shown for SAM-H, and Plittersdorf (right) for all SFMs.
+Right: Rank correlations between SFM IoU and object textural separability, on iShape, Plittersdorf and MOSE for all SFMs.
 6 Robustness to Fine-tuning
-Despite being designed for thin, branching structures, HQ-SAM exhibits performance drops on high tree-likeness objects, similar to SAM and SAM 2 (Table 2, Fig.
+Despite being designed for thin, branching structures, HQ-SAM exhibits performance drops on high tree-likeness objects, similar to SAM and SAM 2 (Fig.
 5).
@@
 To assess whether fine-tuning could improve SAM’s segmentation of high tree-likeness objects, we selected some $k$ examples from DIS-train using different ranking strategies: Oracle (direct IoU-based selection), Rand (random selection), DoGD / CPR (selecting examples based on high tree-likeness), and Sum (a combination of DoGD and CPR).
-We fine-tuned only SAM’s mask decoder on these, keeping the image encoder frozen to preserve its generalist capabilities (as in [24]); full details are in App.
+We fine-tuned only SAM’s mask decoder on these, keeping the image encoder frozen to preserve its generalist capabilities (as in [hqsam]); full details are in App.
 C.3.
@@
 We find that SFMs struggle primarily with over-segmentation rather than under-segmentation in high tree-likeness cases (Sec.
-4.3).
-The models often generate large false-positive regions within densely packed, thin structures (e.g., Figs.
+4.3), generating large false-positive regions within densely packed, thin structures (e.g., Figs.
 3, right; 1, left).
-This cannot be explained by resolution limitations alone, as sparsely packed thin structures exhibit reduced over-segmentation (Fig.
+This cannot be explained by resolution limitations alone, as sparse thin structures exhibit reduced over-segmentation (Fig.
 3, center rightmost).
-Instead, we hypothesize that SFMs misinterpret dense, irregular patterns as texture, confusing these structures with the appearance of a more uniformly shaped object.
-In the cases of both tree-like and low-textural contrast objects, SFMs struggle to delineate object boundaries when faced with misleading textural cues, be it either a dense, irregular shape that appears to instead be a texture (the former case), or an object texture that is similar to its surroundings (the latter).
-This suggests that SAM’s training set (SA-1B) contained few such objects, as tree-like structures and low-contrast objects are uncommon in typical photographic datasets [26] (for the former, see e.g., the typically low object concavity of SA-1B in Fig.
-6, right in [26]).
-Since HQ-SAM and SAM 2 were derived from SAM/its training set, their similar behavior is unsurprising.
+Instead, we hypothesize that SFMs misinterpret dense, irregular patterns as texture, confusing these structures with the appearance of a more uniformly-shaped object.
+For both tree-like and low-contrast objects, SFMs struggle to delineate object boundaries when faced with misleading textural cues, be it either a dense, irregular shape that appears to instead be a texture (the former case), or an object texture similar to its surroundings (the latter).
+This suggests that SAM’s training set (SA-1B) contained few such objects, as tree-like structures and low-contrast objects are uncommon in typical photographic datasets [sam] (for the former, see e.g., the typically low object concavity of SA-1B in Fig.
+6, right in [sam]).
+Since HQ-SAM and SAM 2 were derived from SAM, their similar behavior is unsurprising.
+Comparison to Non-SFM Models.
+Additional experiments with models trained from scratch (IS-Net, iShapeInst, Mask R-CNN) and evaluated on the DIS, iShape and Plittersdorf datasets, respectively (App.
+D.6), revealed that their performance did not sharply degrade on objects with high tree-likeness or low textural separability.
+In contrast to SFMs, such objects were not consistent “failure modes” for these models, suggesting that these difficulties are specific to SFMs rather than being inherently hard cases.
 On HQ-SAM and Fine-tuning.
-Despite being designed and adapted for thin, intricate structures [24], HQ-SAM still exhibits performance drops correlated with tree-likeness (Tables 1, 2), albeit with slight improvements over SAM and SAM 2 (Fig.
+Despite being designed and adapted for thin, intricate structures [hqsam], HQ-SAM still exhibits performance drops correlated with tree-likeness (Figs.
+4, 5), albeit with slight improvements over SAM and SAM 2 (Fig.
 5, left).
@@
 SAM’s image encoder certainly plays a central role, as it serves as the primary feature extractor for the lightweight mask decoder, but a more precise mechanistic analysis could guide architectural improvements to better handle complex structures.
-Conclusion.
-In this paper, we quantitatively modeled how SFM segmentation performance relates to tree-likeness and textural separability.
-Our findings reveal a consistent correlation between these object characteristics and segmentation accuracy, highlighting the need for further research to better understand and mitigate these failure modes—especially since even targeted fine-tuning fails to fully resolve them.
 Acknowledgements
@@
 The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
-Supplementary Material
 Appendix A Additional Dataset Details
@@
 The retinal blood vessel and road satellite image object masks which we use to generate the synthetic images for the experiments in Sec.
-4.3.1, are from the Retina Blood Vessel [22] and the Road Extraction Challenge data of DeepGlobe18 [8].
+4.3.1, are from the Retina Blood Vessel [noauthor_retina_nodate] and the Road Extraction Challenge data of DeepGlobe18 [DeepGlobe18].
 For our retinal vessel mask set, we use the vessel masks from the training set of 80 image/mask pairs, and for our road mask set, we randomly sample the same number of masks from the split-merged full dataset, in order to ensure that these two object types appear equally in the synthetic dataset of Sec.
@@
   1. 1.
-    Sample object mask $m$ from either retinal blood vessel or satellite road image datasets (details in Appendix A.1).
+    Sample object mask $m$ from either retinal blood vessel or satellite road image datasets (details in App. A.1).
   2. 2.
@@
   4. 4.
-    Apply two different textures randomly sampled from Colored Brodatz [1] to the foreground ($m_{c}=1$) and background ($m_{c}=0$), resulting in the final image.
+    Apply two different textures randomly sampled from Colored Brodatz [abdelmounaime2013newMBT] to the foreground ($m_{c}=1$) and background ($m_{c}=0$), resulting in the final image.
 In Fig. 12 we show example images generated with different foreground and background texture combinations for various objects.
@@
 B.1.1 Synthetic Data Experiments
-Tables 4 and 5 show results on the synthetic tree-like dataset using a wide range of CPR and DoGD hyperparameters ($R$ and $a,b$, respectively).
+Tables 1 and 2 show results on the synthetic tree-like dataset using a wide range of CPR and DoGD hyperparameters ($R$ and $a,b$, respectively).
 ViT-H ViT-B
 $R$ $\tau$ $\rho$ $\tau$ $\rho$
-$1$ $-0.82$ $-0.96$ $-0.79$ $-0.94$
-$3$ $-0.80$ $-0.95$ $-0.76$ $-0.94$
-$5$ $-0.77$ $-0.93$ $-0.75$ $-0.93$
-$7$ $-0.74$ $-0.92$ $-0.73$ $-0.91$
-$9$ $-0.72$ $-0.91$ $-0.70$ $-0.90$
-$11$ $-0.71$ $-0.90$ $-0.70$ $-0.89$
-Table 4: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by CPR with different hyperparameter values $R$ on the synthetic dataset, to supplement Table 1.
+11 0.82-0.82 0.96-0.96 0.79-0.79 0.94-0.94
+33 0.80-0.80 0.95-0.95 0.76-0.76 0.94-0.94
+55 0.77-0.77 0.93-0.93 0.75-0.75 0.93-0.93
+77 0.74-0.74 0.92-0.92 0.73-0.73 0.91-0.91
+99 0.72-0.72 0.91-0.91 0.70-0.70 0.90-0.90
+$11$ 0.71-0.71 0.90-0.90 0.70-0.70 0.89-0.89
+Table 1: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by CPR with different hyperparameter values $R$ on the synthetic dataset, to supplement Fig.
+4.
 ViT-H ViT-B
 $a$ $b$ $\tau$ $\rho$ $\tau$ $\rho$
-$63$ $3$ $0.42$ $0.60$ $0.48$ $0.66$
-$63$ $7$ $0.45$ $0.63$ $0.49$ $0.68$
+$63$ 33 $0.42$ $0.60$ $0.48$ $0.66$
+$63$ 77 $0.45$ $0.63$ $0.49$ $0.68$
 $63$ $15$ $0.44$ $0.64$ $0.48$ $0.68$
 $63$ $31$ $0.45$ $0.66$ $0.49$ $0.71$
-$127$ $3$ $0.61$ $0.80$ $0.66$ $0.84$
-$127$ $7$ $0.63$ $0.81$ $0.67$ $0.84$
+$127$ 33 $0.61$ $0.80$ $0.66$ $0.84$
+$127$ 77 $0.63$ $0.81$ $0.67$ $0.84$
 $127$ $15$ $0.62$ $0.81$ $0.66$ $0.84$
 $127$ $31$ $0.60$ $0.80$ $0.64$ $0.83$
-$255$ $3$ $0.66$ $0.84$ $0.68$ $0.85$
-$255$ $7$ $0.66$ $0.84$ $0.68$ $0.86$
+$255$ 33 $0.66$ $0.84$ $0.68$ $0.85$
+$255$ 77 $0.66$ $0.84$ $0.68$ $0.86$
 $255$ $15$ $0.65$ $0.84$ $0.67$ $0.85$
 $255$ $31$ $0.59$ $0.79$ $0.60$ $0.80$
-Table 5: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by DoGD with different hyperparameter values $a,b$ on the synthetic dataset, to supplement Table 1.
+Table 2: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by DoGD with different hyperparameter values $a,b$ on the synthetic dataset, to supplement Fig.
+4.
 B.1.2 Real Data Experiments
-Tables 6, 7, 8, and 9 show results on DIS and iShape using a wide range of CPR and DoGD hyperparameters ($R$ and $a,b$, respectively).
+Tables 3, 4, 5, and 6 show results on DIS and iShape using a wide range of CPR and DoGD hyperparameters ($R$ and $a,b$, respectively).
 ViT-H ViT-B
 $R$ $\tau$ $\rho$ $\tau$ $\rho$
-$1$ $-0.56$ $-0.73$ $-0.61$ $-0.79$
-$3$ $-0.55$ $-0.72$ $-0.61$ $-0.79$
-$5$ $-0.59$ $-0.76$ $-0.63$ $-0.81$
-$7$ $-0.53$ $-0.7$ $-0.59$ $-0.77$
-$9$ $-0.53$ $-0.71$ $-0.57$ $-0.75$
-$11$ $-0.51$ $-0.69$ $-0.56$ $-0.74$
-Table 6: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by CPR with different hyperparameter values $R$ on DIS, to supplement Table 2.
+11 0.56-0.56 0.73-0.73 0.61-0.61 0.79-0.79
+33 0.55-0.55 0.72-0.72 0.61-0.61 0.79-0.79
+55 0.59-0.59 0.76-0.76 0.63-0.63 0.81-0.81
+77 0.53-0.53 0.7-0.7 0.59-0.59 0.77-0.77
+99 0.53-0.53 0.71-0.71 0.57-0.57 0.75-0.75
+$11$ 0.51-0.51 0.69-0.69 0.56-0.56 0.74-0.74
+Table 3: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by CPR with different hyperparameter values $R$ on DIS, to supplement Fig.
+5.
 antenna branch fence hanger log wire
 $R$ SAM Enc. $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
-1 ViT-H $-0.18$ $-0.25$ $-0.63$ $-0.83$ $-0.68$ $-0.88$ $-0.30$ $-0.43$ $-0.60$ $-0.78$ $-0.08$ $-0.12$
-ViT-B $-0.19$ $-0.29$ $-0.64$ $-0.83$ $-0.68$ $-0.88$ $-0.47$ $-0.66$ $-0.61$ $-0.79$ $-0.12$ $-0.18$
-3 ViT-H $-0.20$ $-0.30$ $-0.63$ $-0.83$ $-0.65$ $-0.85$ $-0.28$ $-0.40$ $-0.61$ $-0.79$ $-0.07$ $-0.10$
-ViT-B $-0.17$ $-0.26$ $-0.64$ $-0.83$ $-0.65$ $-0.85$ $-0.48$ $-0.66$ $-0.61$ $-0.79$ $-0.13$ $-0.19$
-5 ViT-H $-0.13$ $-0.21$ $-0.61$ $-0.81$ $-0.61$ $-0.81$ $-0.25$ $-0.37$ $-0.60$ $-0.79$ $-0.06$ $-0.10$
-ViT-B $-0.18$ $-0.27$ $-0.63$ $-0.82$ $-0.63$ $-0.83$ $-0.44$ $-0.62$ $-0.62$ $-0.80$ $-0.12$ $-0.17$
-7 ViT-H $-0.16$ $-0.23$ $-0.57$ $-0.77$ $-0.54$ $-0.74$ $-0.20$ $-0.29$ $-0.60$ $-0.79$ $-0.06$ $-0.10$
-ViT-B $-0.15$ $-0.24$ $-0.59$ $-0.77$ $-0.55$ $-0.75$ $-0.39$ $-0.55$ $-0.62$ $-0.81$ $-0.10$ $-0.15$
-9 ViT-H $-0.16$ $-0.23$ $-0.52$ $-0.72$ $-0.47$ $-0.65$ $-0.13$ $-0.19$ $-0.59$ $-0.78$ $-0.05$ $-0.08$
-ViT-B $-0.16$ $-0.22$ $-0.57$ $-0.76$ $-0.49$ $-0.67$ $-0.29$ $-0.42$ $-0.62$ $-0.81$ $-0.10$ $-0.15$
-11 ViT-H $-0.08$ $-0.13$ $-0.49$ $-0.67$ $-0.51$ $-0.70$ $-0.02$ $-0.03$ $-0.60$ $-0.78$ $-0.03$ $-0.04$
-ViT-B $-0.12$ $-0.16$ $-0.54$ $-0.73$ $-0.55$ $-0.75$ $-0.21$ $-0.28$ $-0.64$ $-0.82$ $-0.07$ $-0.11$
-Table 7: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by CPR with different hyperparameter values $R$ on iShape, to supplement Table 2.
+1 ViT-H 0.18-0.18 0.25-0.25 0.63-0.63 0.83-0.83 0.68-0.68 0.88-0.88 0.30-0.30 0.43-0.43 0.60-0.60 0.78-0.78 0.08-0.08 0.12-0.12
+ViT-B 0.19-0.19 0.29-0.29 0.64-0.64 0.83-0.83 0.68-0.68 0.88-0.88 0.47-0.47 0.66-0.66 0.61-0.61 0.79-0.79 0.12-0.12 0.18-0.18
+3 ViT-H 0.20-0.20 0.30-0.30 0.63-0.63 0.83-0.83 0.65-0.65 0.85-0.85 0.28-0.28 0.40-0.40 0.61-0.61 0.79-0.79 0.07-0.07 0.10-0.10
+ViT-B 0.17-0.17 0.26-0.26 0.64-0.64 0.83-0.83 0.65-0.65 0.85-0.85 0.48-0.48 0.66-0.66 0.61-0.61 0.79-0.79 0.13-0.13 0.19-0.19
+5 ViT-H 0.13-0.13 0.21-0.21 0.61-0.61 0.81-0.81 0.61-0.61 0.81-0.81 0.25-0.25 0.37-0.37 0.60-0.60 0.79-0.79 0.06-0.06 0.10-0.10
+ViT-B 0.18-0.18 0.27-0.27 0.63-0.63 0.82-0.82 0.63-0.63 0.83-0.83 0.44-0.44 0.62-0.62 0.62-0.62 0.80-0.80 0.12-0.12 0.17-0.17
+7 ViT-H 0.16-0.16 0.23-0.23 0.57-0.57 0.77-0.77 0.54-0.54 0.74-0.74 0.20-0.20 0.29-0.29 0.60-0.60 0.79-0.79 0.06-0.06 0.10-0.10
+ViT-B 0.15-0.15 0.24-0.24 0.59-0.59 0.77-0.77 0.55-0.55 0.75-0.75 0.39-0.39 0.55-0.55 0.62-0.62 0.81-0.81 0.10-0.10 0.15-0.15
+9 ViT-H 0.16-0.16 0.23-0.23 0.52-0.52 0.72-0.72 0.47-0.47 0.65-0.65 0.13-0.13 0.19-0.19 0.59-0.59 0.78-0.78 0.05-0.05 0.08-0.08
+ViT-B 0.16-0.16 0.22-0.22 0.57-0.57 0.76-0.76 0.49-0.49 0.67-0.67 0.29-0.29 0.42-0.42 0.62-0.62 0.81-0.81 0.10-0.10 0.15-0.15
+11 ViT-H 0.08-0.08 0.13-0.13 0.49-0.49 0.67-0.67 0.51-0.51 0.70-0.70 0.02-0.02 0.03-0.03 0.60-0.60 0.78-0.78 0.03-0.03 0.04-0.04
+ViT-B 0.12-0.12 0.16-0.16 0.54-0.54 0.73-0.73 0.55-0.55 0.75-0.75 0.21-0.21 0.28-0.28 0.64-0.64 0.82-0.82 0.07-0.07 0.11-0.11
+Table 4: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by CPR with different hyperparameter values $R$ on iShape, to supplement Fig.
+5.
 ViT-H ViT-B
 $a$ $b$ $\tau$ $\rho$ $\tau$ $\rho$
-$63$ $3$ $0.63$ $0.83$ $0.57$ $0.78$
-$63$ $7$ $0.59$ $0.79$ $0.57$ $0.77$
+$63$ 33 $0.63$ $0.83$ $0.57$ $0.78$
+$63$ 77 $0.59$ $0.79$ $0.57$ $0.77$
 $63$ $15$ $0.54$ $0.72$ $0.50$ $0.67$
 $63$ $31$ $0.45$ $0.62$ $0.39$ $0.56$
-$127$ $3$ $0.58$ $0.77$ $0.45$ $0.64$
-$127$ $7$ $0.58$ $0.77$ $0.49$ $0.66$
+$127$ 33 $0.58$ $0.77$ $0.45$ $0.64$
+$127$ 77 $0.58$ $0.77$ $0.49$ $0.66$
 $127$ $15$ $0.51$ $0.67$ $0.42$ $0.57$
 $127$ $31$ $0.36$ $0.50$ $0.26$ $0.39$
-$255$ $3$ $0.53$ $0.71$ $0.42$ $0.59$
-$255$ $7$ $0.53$ $0.71$ $0.42$ $0.59$
+$255$ 33 $0.53$ $0.71$ $0.42$ $0.59$
+$255$ 77 $0.53$ $0.71$ $0.42$ $0.59$
 $255$ $15$ $0.47$ $0.66$ $0.36$ $0.52$
 $255$ $31$ $0.29$ $0.43$ $0.19$ $0.29$
-Table 8: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by DoGD with different hyperparameter values $a,b$ on DIS, to supplement Table 2.
+Table 5: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by DoGD with different hyperparameter values $a,b$ on DIS, to supplement Fig.
+5.
 antenna branch fence hanger log wire
 $a$ $b$ SAM Enc. $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
-63 3 ViT-H $0.39$ $0.55$ $0.62$ $0.82$ $0.65$ $0.85$ $-0.15$ $-0.23$ $0.23$ $0.35$ $0.35$ $0.50$
-ViT-B $0.25$ $0.35$ $0.62$ $0.81$ $0.60$ $0.81$ $-0.07$ $-0.11$ $0.17$ $0.25$ $0.36$ $0.51$
-63 7 ViT-H $0.28$ $0.38$ $0.61$ $0.81$ $0.62$ $0.83$ $-0.13$ $-0.21$ $0.22$ $0.33$ $0.34$ $0.49$
-ViT-B $0.13$ $0.18$ $0.61$ $0.81$ $0.60$ $0.80$ $-0.07$ $-0.11$ $0.14$ $0.22$ $0.33$ $0.48$
-63 15 ViT-H $0.10$ $0.12$ $0.59$ $0.78$ $0.59$ $0.80$ $-0.05$ $-0.07$ $0.19$ $0.30$ $0.29$ $0.43$
-ViT-B $0.04$ $0.03$ $0.59$ $0.77$ $0.56$ $0.77$ $-0.01$ $-0.01$ $0.09$ $0.16$ $0.25$ $0.36$
+63 3 ViT-H $0.39$ $0.55$ $0.62$ $0.82$ $0.65$ $0.85$ 0.15-0.15 0.23-0.23 $0.23$ $0.35$ $0.35$ $0.50$
+ViT-B $0.25$ $0.35$ $0.62$ $0.81$ $0.60$ $0.81$ 0.07-0.07 0.11-0.11 $0.17$ $0.25$ $0.36$ $0.51$
+63 7 ViT-H $0.28$ $0.38$ $0.61$ $0.81$ $0.62$ $0.83$ 0.13-0.13 0.21-0.21 $0.22$ $0.33$ $0.34$ $0.49$
+ViT-B $0.13$ $0.18$ $0.61$ $0.81$ $0.60$ $0.80$ 0.07-0.07 0.11-0.11 $0.14$ $0.22$ $0.33$ $0.48$
+63 15 ViT-H $0.10$ $0.12$ $0.59$ $0.78$ $0.59$ $0.80$ 0.05-0.05 0.07-0.07 $0.19$ $0.30$ $0.29$ $0.43$
+ViT-B $0.04$ $0.03$ $0.59$ $0.77$ $0.56$ $0.77$ 0.01-0.01 0.01-0.01 $0.09$ $0.16$ $0.25$ $0.36$
 63 31 ViT-H $0.06$ $0.04$ $0.55$ $0.74$ $0.46$ $0.65$ $0.07$ $0.10$ $0.14$ $0.26$ $0.09$ $0.13$
-ViT-B $-0.11$ $-0.17$ $0.56$ $0.74$ $0.45$ $0.64$ $0.07$ $0.11$ $0.02$ $0.08$ $0.02$ $0.03$
+ViT-B 0.11-0.11 0.17-0.17 $0.56$ $0.74$ $0.45$ $0.64$ $0.07$ $0.11$ $0.02$ $0.08$ $0.02$ $0.03$
 127 3 ViT-H $0.23$ $0.30$ $0.52$ $0.70$ $0.46$ $0.65$ $0.52$ $0.72$ $0.30$ $0.46$ $0.08$ $0.11$
 ViT-B $0.12$ $0.14$ $0.51$ $0.69$ $0.45$ $0.64$ $0.50$ $0.70$ $0.13$ $0.22$ $0.06$ $0.08$
-127 7 ViT-H $0.23$ $0.26$ $0.48$ $0.65$ $0.41$ $0.59$ $0.51$ $0.70$ $0.29$ $0.46$ $-0.01$ $-0.01$
-ViT-B $0.10$ $0.09$ $0.51$ $0.68$ $0.41$ $0.59$ $0.49$ $0.68$ $0.13$ $0.22$ $-0.03$ $-0.05$
-$127$ $15$ ViT-H $0.14$ $0.17$ $0.43$ $0.59$ $0.38$ $0.55$ $0.50$ $0.69$ $0.28$ $0.43$ $-0.09$ $-0.12$
-ViT-B $0.07$ $0.03$ $0.45$ $0.62$ $0.36$ $0.52$ $0.48$ $0.67$ $0.12$ $0.20$ $-0.14$ $-0.20$
-$127$ $31$ ViT-H $0.05$ $0.05$ $0.36$ $0.50$ $0.08$ $0.12$ $0.48$ $0.67$ $0.30$ $0.45$ $-0.15$ $-0.22$
-ViT-B $0.04$ $0.00$ $0.38$ $0.52$ $0.12$ $0.18$ $0.46$ $0.65$ $0.12$ $0.20$ $-0.21$ $-0.31$
-$255$ $3$ ViT-H $0.04$ $0.06$ $0.33$ $0.46$ $-0.20$ $-0.30$ $0.44$ $0.61$ $0.52$ $0.72$ $-0.17$ $-0.25$
-ViT-B $0.21$ $0.31$ $0.30$ $0.42$ $-0.13$ $-0.20$ $0.42$ $0.59$ $0.42$ $0.60$ $-0.23$ $-0.34$
-$255$ $7$ ViT-H $0.02$ $0.02$ $0.29$ $0.41$ $-0.25$ $-0.35$ $0.43$ $0.60$ $0.51$ $0.71$ $-0.20$ $-0.29$
-ViT-B $0.19$ $0.27$ $0.26$ $0.36$ $-0.18$ $-0.26$ $0.42$ $0.58$ $0.41$ $0.59$ $-0.25$ $-0.36$
-$255$ $15$ ViT-H $-0.04$ $-0.06$ $0.24$ $0.34$ $-0.31$ $-0.45$ $0.43$ $0.61$ $0.50$ $0.70$ $-0.20$ $-0.30$
-ViT-B $0.21$ $0.29$ $0.20$ $0.28$ $-0.26$ $-0.39$ $0.40$ $0.56$ $0.40$ $0.58$ $-0.28$ $-0.40$
-$255$ $31$ ViT-H $-0.11$ $-0.14$ $0.11$ $0.16$ $-0.47$ $-0.65$ $0.43$ $0.59$ $0.46$ $0.66$ $-0.23$ $-0.33$
-ViT-B $0.13$ $0.20$ $0.07$ $0.09$ $-0.42$ $-0.60$ $0.39$ $0.55$ $0.38$ $0.55$ $-0.30$ $-0.44$
-Table 9: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by DoGD with different hyperparameter values $R$ on iShape, to supplement Table 2.
+127 7 ViT-H $0.23$ $0.26$ $0.48$ $0.65$ $0.41$ $0.59$ $0.51$ $0.70$ $0.29$ $0.46$ 0.01-0.01 0.01-0.01
+ViT-B $0.10$ $0.09$ $0.51$ $0.68$ $0.41$ $0.59$ $0.49$ $0.68$ $0.13$ $0.22$ 0.03-0.03 0.05-0.05
+$127$ $15$ ViT-H $0.14$ $0.17$ $0.43$ $0.59$ $0.38$ $0.55$ $0.50$ $0.69$ $0.28$ $0.43$ 0.09-0.09 0.12-0.12
+ViT-B $0.07$ $0.03$ $0.45$ $0.62$ $0.36$ $0.52$ $0.48$ $0.67$ $0.12$ $0.20$ 0.14-0.14 0.20-0.20
+$127$ $31$ ViT-H $0.05$ $0.05$ $0.36$ $0.50$ $0.08$ $0.12$ $0.48$ $0.67$ $0.30$ $0.45$ 0.15-0.15 0.22-0.22
+ViT-B $0.04$ $0.00$ $0.38$ $0.52$ $0.12$ $0.18$ $0.46$ $0.65$ $0.12$ $0.20$ 0.21-0.21 0.31-0.31
+$255$ 33 ViT-H $0.04$ $0.06$ $0.33$ $0.46$ 0.20-0.20 0.30-0.30 $0.44$ $0.61$ $0.52$ $0.72$ 0.17-0.17 0.25-0.25
+ViT-B $0.21$ $0.31$ $0.30$ $0.42$ 0.13-0.13 0.20-0.20 $0.42$ $0.59$ $0.42$ $0.60$ 0.23-0.23 0.34-0.34
+$255$ 77 ViT-H $0.02$ $0.02$ $0.29$ $0.41$ 0.25-0.25 0.35-0.35 $0.43$ $0.60$ $0.51$ $0.71$ 0.20-0.20 0.29-0.29
+ViT-B $0.19$ $0.27$ $0.26$ $0.36$ 0.18-0.18 0.26-0.26 $0.42$ $0.58$ $0.41$ $0.59$ 0.25-0.25 0.36-0.36
+$255$ $15$ ViT-H 0.04-0.04 0.06-0.06 $0.24$ $0.34$ 0.31-0.31 0.45-0.45 $0.43$ $0.61$ $0.50$ $0.70$ 0.20-0.20 0.30-0.30
+ViT-B $0.21$ $0.29$ $0.20$ $0.28$ 0.26-0.26 0.39-0.39 $0.40$ $0.56$ $0.40$ $0.58$ 0.28-0.28 0.40-0.40
+$255$ $31$ ViT-H 0.11-0.11 0.14-0.14 $0.11$ $0.16$ 0.47-0.47 0.65-0.65 $0.43$ $0.59$ $0.46$ $0.66$ 0.23-0.23 0.33-0.33
+ViT-B $0.13$ $0.20$ $0.07$ $0.09$ 0.42-0.42 0.60-0.60 $0.39$ $0.55$ $0.38$ $0.55$ 0.30-0.30 0.44-0.44
+Table 6: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object tree-likeness as measured by DoGD with different hyperparameter values $R$ on iShape, to supplement Fig.
+5.
 B.2 Object Textural Separability Experiments
@@
 ViT-B $0.58$ $0.77$ $0.52$ $0.70$ $0.54$ $0.74$ $0.54$ $0.74$ $0.23$ $0.33$ $0.41$ $0.57$ $0.15$ $0.23$
-Table 10: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object textural separability (Algorithm 3), on real datasets (iShape and Plittersdorf), using different weak classifier models $g$ and hyperparameters to measure textural separability (to supplement Table 3).
+Table 7: Rank correlation (Kendall $\tau$ and Spearman $\rho$) between SAM prediction IoU and object textural separability (Algorithm 3), on real datasets (iShape and Plittersdorf), using different weak classifier models $g$ and hyperparameters to measure textural separability (to supplement Fig.
+9).
 Appendix C Additional Experimental Details
@@
 3, all real and style-transferred (NST) images are prompted with (1) a tight bounding-box and (2) several positive and/or negative prompts randomly sampled from either the foreground or background of the object mask, respectively, with respective counts $n_{pos}$ and $n_{neg}$.
-For iShape, we use $n_{pos}=n_{neg}=5$; for DIS, we use $n_{pos}=5$ and $n_{neg}=10$ due to the complexity of the objects.
+For iShape, we use $n_{pos}=n_{neg}=5$; for DIS and MOSE, we use $n_{pos}=5$ and $n_{neg}=10$ due to the complexity of the objects.
 For Plittersdorf and NST images, we simply use $n_{pos}=0$ and $n_{neg}=2$, due to the objects typically possessing simple shapes.
@@
 5.2.1), the inpainting model used is Runway’s stable diffusion inpainting pipeline, fp16 variant.
-The NST model itself is an implementation of [14] based on https://github.com/pytorch/tutorials/blob/main/advanced_source/neural_style_tutorial.py, using typical settings of a standard VGG19 CNN with content layer of conv_4, style layers of conv_i for i = 1, 2, 3, 4, 5, and ImageNet normalization.
+The NST model itself is an implementation of [gatys2016image] based on https://github.com/pytorch/tutorials/blob/main/advanced_source/neural_style_tutorial.py, using typical settings of a standard VGG19 CNN with content layer of conv_4, style layers of conv_i for i = 1, 2, 3, 4, 5, and ImageNet normalization.
 Style transfer intensity.
@@
   1. 1.
-    Apply a non-affine transformation created using the albumentations Python library [4] to both the image and the object/mask, defined shortly.
+    Apply a non-affine transformation created using the albumentations Python library [info11020125] to both the image and the object/mask, defined shortly.
   2. 2.
@@
 Here we provide all details for the SAM fine-tuning experiments of Section 6.
-The loss function that we use is the sum of the Dice and cross-entropy losses (as implemented by MONAI [5]), optimized via AdamW [29] with a learning rate of $10^{-5}$ and a weight decay strength of $0.1$.
+The loss function that we use is the sum of the Dice and cross-entropy losses (as implemented by MONAI [cardoso2022monai]), optimized via AdamW [loshchilov2018decoupled] with a learning rate of $10^{-5}$ and a weight decay strength of $0.1$.
 We use a batch size of $16$ and fine-tune for $40$ epochs on the $k$ (image, mask) examples which were sampled from DIS-train using the given selection strategy (see Section 6).
@@
 Appendix D Additional Results
-D.1 Correlation Between Metrics
+D.1 SAM Self-Attention Maps
 In Fig.
-D.1, we show the relationship between our proposed tree-likeness metrics CPR and DoGD on all three datasets (with default hyperparameters and ViT-H SAM), quantified by correlations shown in Table 11.
+14 we show examples of SAM’s attention patterns for images with objects of high tree-likeness.
+These were computed via ground truth mask-guided attention rollout [abnar-zuidema-2020-quantifying].
+Figure 14: Attention patterns (right column) for example images (left column) with objects of high tree-likeness (shown via ground truth masks in middle column).
+D.2 Mechanistic Relationships between Metrics and Model Architectural Choices
+D.2.1 The Relationship of Vision Transformer Effective Patch Size with CPR
+As was described in Sec.
+4.1, in theory, when the patch size of an SFM’s image encoder is small relative to the typical branch width of a tree-like object, patches can still contain internal pixels, but as patch size grows, CPR implies that boundary pixels dominate every patch, leaving no truly “interior” representation.
+Thus, CPR should directly interact with patch size: larger patches exacerbate the imbalance between boundary and interior content, i.e., the typical CPR of objects will be higher.
+In this section, we will verify this behavior experimentally.
+As the patch sizes of the SFMs (SAM, HQ-SAM and SAM 2) are fixed a priori, we propose to instead modify the effective patch size by altering the size of the objects themselves.
+We do this by zooming a given image in (cropping) our out, centered on the object mask, such that the length of the longer edge of the object’s tight bounding box is of size $p$ pixels (and similarly modifying the object mask accordingly).
+In the case of zooming out, we zero-pad the image to set it to the proper size for SFM input, $1024\times 1024$.
+For example, if $p$ is doubled (the object is made twice as large), the object will be covered by roughly four times as many patches, making the effective patch size halved, because more patches are required to cover the same object.
+In summary, higher $p$ corresponds to smaller effective patch size.
+In Fig.
+15, we show how the performance (IoU) vs.
+object CPR curve on DIS is affected by different choices of $p\in[32,64,96,160,224,352,480,736]$, for both variants of SAM.
+We note that we discard any datapoint where the zooming-in operation made the object grow outside of the image boundaries.
+We see that as was predicted, larger patch sizes (smaller $p$) results in objects typically having noticeably higher CPR (higher average CPR $\mu$ for a given $p$, as shown in the plot legends).
+Moreover, the higher typical CPR results in the model performance on a given object being typically worse.
+Figure 15: The effect on the relationship between segmentation performance (IoU) vs.
+CPR (object tree-likeness) of changing the model’s effective patch size, increased with decreasing $p$, computed on DIS for the SAM ViT-H (left) and ViT-B (right) variants.
+Also shown are each plot’s legend are average CPR values $\mu$ and estimated linear slope of IoU vs.
+CPR, for each pp.
+A potential mechanistic solution to this problem could be to use vision transformer models with smaller patch sizes to perform better on tree-like objects, yet this carries a high computational burden, because the size of the self-attention matrices grows quadratically with the patch size becoming smaller.
+D.2.2 The Relationship of Vision Transformer Attention Map Fragmentation with CPR
+As we mentioned in Section 4.1 and showed via a few examples in App.
+D.1, attention maps on tree-like objects typically appear to be fragmented and locally-focused, due to responding to local edges independently instead of forming clear global object representations.
+Here, we will test this empirically by determining if the correlation of the fragmentation of an input image’s attention map with the tree-likeness of the object that the image contains.
+Here, the attention map for a given input image is defined via the attention rollout technique used in App.
+D.1.
+We measure the fragmentation of a self-attention map $A\in[0,1]^{K\times K}$ via Moran’s spatial autocorrelation index $I$ [moran1950notes].
+Moran’s $I$ is lower for more fragmented/less clustered arrays, so we would expect theoretically that the CPR/tree-likeness of an object’s mask is negatively correlated with the Moran’s $I$ of the model’s computed attention map for the image containing the object.
+To empirically measure the relationship between an SFM attention map’s fragmentation and the tree-likeness of an object to be segmented, we measure the correlation of the Moran’s $I$ of an input image’s self attention map with the CPR of the input image’s associated mask label, computed for both SAM-H and SAM-B on DIS and all classes of iShape, using the same aggregation procedure as in the main experiments (Sec.
+4.3.2).
+The results are shown in Table 8, with accompanying plots in Fig.
+16.
+Overall, we indeed see a strong negative correlation between object CPR and self-attention map Moran’s $I$: objects of higher tree-likeness correspond to more fragmented self-attention maps, as was indicated by our qualitative results in App.
+D.1.
+SFM Model DIS iShape (all classes)
+$\tau$ $\rho$ $\tau$ $\rho$
+SAM-H -0.72 -0.89 -0.76 -0.91
+SAM-B -0.59 -0.78 -0.72 -0.88
+Table 8: Nonlinear (Spearman $\rho$, Kendall $\tau$) correlations of self-attention map fragmentation (Moran’s autocorrelation II) with object tree-likeness (CPR).
+Figure 16: Object tree-likeness (CPR) vs.
+self-attention map fragmentation (Moran’s autocorrelation II), on DIS (top row) and iShape (bottom row) for both ViT-H (left column) and ViT-B (right column) variants of SAM.
+D.3 Correlation Between Metrics
+In Fig.
+D.3, we show the relationship between our proposed tree-likeness metrics CPR and DoGD on all three datasets (with default hyperparameters and ViT-H SAM), quantified by correlations shown in Table 9.
 Figure 17: Relationship between tree-likeness metrics CPR and DoGD. From left to right, on synthetic data (Fig. 4), DIS, and iShape (Fig. 5).
@@
 Synth. DIS antenna branch fence hanger log wire all Avg.
-Pearson $r$ $-0.94$ $-0.67$ $-0.92$ $-0.93$ $-0.82$ $-0.81$ $-0.90$ $-0.66$ $-0.79$ $-0.83$
-Spearman $\rho$ $-0.82$ $-0.70$ $-0.91$ $-0.93$ $-0.86$ $-0.76$ $-0.94$ $-0.79$ $-0.76$ $-0.83$
-Kendall $\tau$ $-0.62$ $-0.51$ $-0.75$ $-0.78$ $-0.68$ $-0.58$ $-0.79$ $-0.61$ $-0.57$ $-0.65$
+Pearson $r$ 0.94-0.94 0.67-0.67 0.92-0.92 0.93-0.93 0.82-0.82 0.81-0.81 0.90-0.90 0.66-0.66 0.79-0.79 0.83-0.83
+Spearman $\rho$ 0.82-0.82 0.70-0.70 0.91-0.91 0.93-0.93 0.86-0.86 0.76-0.76 0.94-0.94 0.79-0.79 0.76-0.76 0.83-0.83
+Kendall $\tau$ 0.62-0.62 0.51-0.51 0.75-0.75 0.78-0.78 0.68-0.68 0.58-0.58 0.79-0.79 0.61-0.61 0.57-0.57 0.65-0.65
 Table 9: Linear (Pearson $r$) and nonlinear (Spearman $\rho$, Kendall $\tau$) correlations between CPR and DoGD on all evaluated datasets.
-D.2 Effect of Object Line Thickness on Performance
+D.4 Effect of Object Line Thickness on Performance
 In Section 4.1, we hypothesized that the noticeable SAM segmentation performance difference between retinal vessel images and satellite road images, two types of objects with branching structures, is due to the former objects typically having more dense and irregularly-spaced branching structures than the latter.
@@
 We show the results of this procedure (how it affects the relationship between segmentation IoU and tree-likeness) in Fig.
-15.
+18.
 We see that the result of IoU and object tree-likeness being correlated is mostly invariant to the object line thickening transformation.
@@
 Figure 18: SAM segmentation IoU vs. object tree-likeness (CPR, left; DoGD, right) on the synthetic dataset with dilated object lines.
+D.5 Performance vs. NST intensity for all SFMs
+To supplement Fig.
+8 right, we show the same results of segmentation IoU vs.
+NST intensity for all groups (controlled, altered, and mixed) for the additional SFMs (SAM ViT-B, SAM2 ViT-B+ and ViT-L, HQ-SAM ViT-H and ViT-B) in Fig.
+19.
+Figure 19: From left-to-right and top-to-bottom: segmentation IoU vs.
+NST intensity for all groups (controlled, altered, and mixed) for SAM ViT-B, SAM 2 ViT-B+, SAM 2 ViT-L, HQ-SAM ViT-H, and HQ-SAM ViT-B.
+D.6 Comparison to Segmentation Models Trained From Scratch
+We additionally evaluate training models from scratch on each dataset to see if the discovered relationships of performance with tree-likeness (CPR and DoGD) and textural separability metrics is indeed a weakness unique to segmentation foundation models (SFMs), or if such cases are simply universally challenging from a general standpoint, and cannot be solved by training a model from scratch for the specific dataset.
+For DIS, we use the pretrained IS-Net model available directly from the original paper [DIS], obtained from https://github.com/xuebinqin/DIS.
+For iShape, we use the iShapeInst models from wang2025ishapeinst, obtained from https://github.com/youngnuaa/iShapeInst.
+We defer to those works for all architectural and training details.
+For Plittersdorf, no existing pretrained models could be found, so we trained a Mask R-CNN [he2017mask] on the Plittersdorf training set, fine-tuning from COCO weights (we could not train completely from scratch due to the relatively small size of this dataset), and tested on the test set.
+The Plittersdorf model utilized a ResNet-50-FPN backbone, with the COCO-InstanceSegmentation/mask_rcnn_R_50_FPN_3x.yaml configuration of the Detectron2 [wu2019detectron2] model zoo.
+For training, we used a batch size of 2, a learning rate of $2.5\times 10^{-4}$, $30,000$ total training iterations, $300$ warmup iterations, and a learning rate schedule that reduced the learning rate by factor of 0.1 at 60% and 80% of total training time.
+We show all results in Fig.
+20 for DIS, Fig.
+21 and Table 10 for iShape, and Fig.
+22 for Plittersdorf.
+There are several important takeaways when we compare these results of models trained from scratch to the behavior of the segmentation foundation models in the main text.
+First, in all cases, we see that the models perform noticeably better (in average IoU) than the SFMs, which makes sense given that they were trained specifically and only on these datasets, which is unsurprising.
+In the case of object tree-likeness (on DIS and iShape), we see that the sharpness (approximate slope) of the effect of CPR and DoGD on performance is far lower than compared to the SFM results in the main text (Sec.
+4.3.2).
+In other words, these models typically perform much better on objects of high tree-likeness (i.e., low CPR/high DoGD) than the SFMs, indicating that they do not share the same “failure mode” behavior.
+The correlations between IoU and tree-likeness are lower as well for these models trained from scratch (comparing Fig.
+5 for the SFMs to Figs.
+20 and 21 and Table 10 here).
+We see similar, if not stronger results for textural separability: there is, on average, little to no correlation of performance with separability on both iShape (Fig.
+21 and Table 10) and Plittersdorf (Fig.
+22), compared to the noticeably stronger effect of an object’s textural separability on SFM performance (Fig.
+9).
+We hypothesize that any remaining correlation of performance with these object characteristics (tree-likeness and textural separability) is simply because as shown in the figures, objects with higher tree-likeness or lower textural separability are simply less common in the dataset, so models trained from scratch will perform slightly worse on such cases compared to on cases which have more examples to learn from.
+Overall, these results support that SFMs are uniquely susceptible to objects of high tree-likeness or low textural separability compared to models trained from scratch on datasets that have these types of uncommon objects.
+Figure 20: For IS-Net model trained from scratch on DIS: relationship of performance (IoU on the DIS validation set) with tree-likeness metrics: CPR, left and DoGD, right.
+For CPR, Spearman correlation $\rho=-0.61$ and Kendall $\tau=-0.45$.
+For DoGD, $\rho=0.46$ and $\tau=0.32$.
+Figure 21: For iShapeInst models trained from scratch on iShape: relationship of performance (IoU) with tree-likeness metrics of CPR (left) and DoGD (center), and with textural separability (right).
+See Table 10 for numerical correlation results.
+antenna branch fence hanger log wire Avg.
+CPR Spearman $\rho$ -0.31 -0.78 -0.61 -0.36 -0.53 -0.08 -0.45
+Kendall $\tau$ -0.21 -0.59 -0.45 -0.25 -0.38 -0.05 -0.32
+DoGD Spearman $\rho$ 0.08 0.55 0.39 0.16 0.36 -0.01 0.26
+Kendall $\tau$ 0.06 0.39 0.26 0.11 0.25 -0.01 0.18
+Textural
+Separability
+Spearman $\rho$ 0.34 0.40 0.51 0.43 0.10 0.22 0.33
+Kendall $\tau$ 0.23 0.27 0.35 0.30 0.07 0.15 0.23
+Table 10: For iShape, nonlinear (Spearman $\rho$, Kendall $\tau$) correlations between performance of models trained from scratch (IoU) and both tree-likeness metrics (CPR and DoGD) as well as textural separability.
+Figure 22: For Mask R-CNN model trained on Plittersdorf: relationship of performance (IoU) with textural separability (Spearman correlation $\rho=-0.07$ and Kendall $\tau=-0.05$.)
+D.7 All results on MOSE
+In Fig.
+23 we show the corresponding plots for the SFM performance (IoU) vs.
+tree-likeness (table in Fig.
+5) and textural separability (table in Fig.
+9 relationship results on the MOSE dataset.
+These are for the same SFM architectures as shown in the plots of Figs.
+5) and 9).
+Figure 23: Segmentation IoU vs. object tree-likeness (CPR, left, and DoGD, center) and textural separability (right) for the MOSE dataset, for all three evaluated SFMs.